### 24 Mar 2026 v1

# Critical LAN and Score Tests for Mixed Fractional Models under High-Frequency Observation at $H=3/4$

Abstract
We study the critical boundary $H=3/4$ for two mixed fractional models under high-frequency observation, namely mixed fractional Brownian motion and mixed fractional Ornstein–Uhlenbeck. For different reasons, both the supercritical normalization for $H>3/4$ and the subcritical normalization for $H<3/4$ fail at this value. We identify the exact critical first-order scales, prove the corresponding score central limit theorems, and show that once the explicit linear term in the $H$-score is removed, the resulting $(\sigma,H)$-block is already non-degenerate. Thus, in contrast with the regime $H>3/4$, the critical point is resolved by a single triangular local reparametrization. Building on this reduction, we derive the critical second-order likelihood expansion and establish LAN for both models with fully explicit leading information constants.
Motivated by a testing viewpoint from mathematical finance, we additionally derive score-type tests at the critical boundary. In particular, we formulate one-sided procedures for detecting whether the system enters the supercritical side $H>3/4$. The proposed tests are calibrated under the critical constraint $H=3/4$, use a right-tail rejection region dictated by the LAN drift, and admit fully explicit normalizations at the leading order.
1 Introduction
Mixed fractional models play an important role in financial econometrics and the statistics of stochastic processes. By superimposing a standard Brownian motion with a fractional Brownian motion, these models capture long-range dependence while preserving a semimartingale structure, thereby ruling out arbitrage opportunities in frictionless markets [4]. In recent years, high-frequency statistical inference for the mixed fractional Brownian motion (mfBm) and the mixed fractional Ornstein–Uhlenbeck (mfOU) process has attracted considerable attention.
For fractional noise models, the low frequency behaviour determines the asymptotic properties of estimators. [1] established local asymptotic normality (LAN) for fractional Gaussian noise under high-frequency observations, revealing the role of the Hurst exponent $H$. [2] extended the LAN theory to mixed fractional Brownian motion and showed that for $H>3/4$ two non-diagonal transformations are required to regularise the parameter space, whereas for $H<3/4$ a single transformation suffices. [3] obtained analogous results for the mixed fractional OU process and further demonstrated that the $\alpha$-direction is asymptotically orthogonal to the $(\sigma,H)$-block under the critical scaling.
Nevertheless, all the above works exclude the boundary $H=3/4$. At this critical point, the low frequency singularity of the spectral density achieves a precise balance between signal and noise, causing both the supercritical and subcritical normalisations to break down. Specifically, when $H=3/4$ the transformed $(\sigma,H)$-block is no longer singular, but the leading traces acquire an extra logarithmic factor, necessitating a novel scaling. The statistical inference at this critical boundary has remained unexplored in the existing literature.
From a practical perspective, detecting whether the system enters the supercritical region $H>3/4$ is of particular interest in continuous time finance, as it signals stronger long-range dependence and may affect risk management and derivative pricing. Consequently, developing a testing procedure that is calibrated at the boundary $H=3/4$ and has a clear directional interpretation becomes a natural objective.
This paper fills the above gaps. We identify the exact first-order scales at $H=3/4$, prove central limit theorems for the transformed score vectors, and establish a complete LAN expansion whose information matrix is given by fully explicit constants for two models
  1. (i)
    mixed fractional Brownian motion (mfBm),
  2. (ii)
    mixed fractional Ornstein–Uhlenbeck (mfOU).
In many continuous time financial models, semimartingale type restrictions are invoked to rule out arbitrage under frictionless trading. Mixed fractional models exhibit a pronounced change of local asymptotic structure around the boundary $H=3/4$. Thus this boundary is a natural point at which to formulate statistical tests for regime identification.
In this paper, we focus on the one-sided testing problem
$H_{0}:\ H\leq\frac{3}{4}\qquad\text{vs}\qquad H_{1}:\ H>\frac{3}{4},$
which is natural from the supercritical viewpoint. Since the score statistics are constructed at $H=3/4$, the sign of the local LAN drift in the $H$-direction leads to a right-tailed rejection region. For completeness, we also record the corresponding two-sided boundary test.
Notation.
  • For an integrable symbol $g$ on $\pi,\pi]$, let
    $\mathcal{T}_{n}(g):=\bigl(\widehat{g}(i-j)\bigr)_{1\leq i,j\leq n},\qquad\widehat{g}(k):=\frac{1}{2\pi}\int_{-\pi}^{\pi}g(\lambda)e^{-ik\lambda}\,\mathrm{d}\lambda.$
    We reserve $T_{n}=n\Delta_{n}$ for the observation horizon; this removes the ambiguity between the time horizon and the Toeplitz operator. If $A_{n}$ is a real symmetric matrix and $Z_{n}\sim\mathcal{N}(0,I_{n})$, write
    $Q_{n}(A_{n}):=Z_{n}^{\top}A_{n}Z_{n}-\mathrm{tr}(A_{n}).$ (1)
  • Throughout, we assume that
    $\Delta_{n}\downarrow 0,\qquad T_{n}:=n\Delta_{n}\to\infty,\qquad L_{n}:=\log(1/\Delta_{n}).$
    All matrices and constants are evaluated at the critical point $H=3/4$.
The paper is organized as follows. Section 2 introduces the mfBm and mfOU models, derives the critical low frequency constants and covariance matrices at $H=3/4$, and establishes the exact score decompositions and critical score profiles used throughout this paper. Section 3 states the critical score CLTs and the LAN theorems for the mfBm and mfOU models. Sections 4 and 5 prove these results. Section 6 develops score-type tests at the critical boundary $H=3/4$. Section 7 reports a simulation study for the feasible mfBm test. Appendix A collects the low frequency reductions, the fully explicit constants and the higher-order remainder bounds.
2 Preliminaries
In this section, we collect the model setup, notation and critical quantities needed for the analysis at the critical point $H=3/4$. We first introduce the mfBm and mfOU models, together with their low frequency spectral behavior at the critical point. We then define the corresponding critical covariance constants and matrices, derive the exact Gaussian score decompositions, and define the critical score profiles.
2.1 The mfBm model
Let
$Y_{t}=\sigma B_{t}^{H}+W_{t},\qquad t\geq 0,$
where $B_{t}^{H}$ is a fractional Brownian motion, $W_{t}$ is an independent standard Brownian motion. The parameters $(\sigma,H)$ are assumed known. We observe the increment vector
$X_{n}^{\mathrm{m}}:=(\Delta_{1}^{n}Y,\dots,\Delta_{n}^{n}Y)^{\top},\qquad\Delta_{k}^{n}Y:=Y_{k\Delta_{n}}-Y_{(k-1)\Delta_{n}},$
where the superscript $\mathrm{m}$ denotes that the observations are from the mfBm model. Its covariance matrix is
$\Sigma_{n}^{\mathrm{m}}(\sigma,H)=\Delta_{n}\Bigl(I_{n}+\sigma^{2}\Delta_{n}^{2H-1}\mathcal{T}_{n}(f_{H})\Bigr),$
where $f_{H}$ is the spectral density of fractional Gaussian noise. Under the Fourier convention fixed in notation, one may take
$f_{H}(\lambda)=2\,\Gamma(2H+1)\sin(\pi H)\,(1-\cos\lambda)\sum_{k\in\mathbb{Z}}\left|\lambda+2\pi k\right|^{-(2H+1)},\qquad\lambda\in[-\pi,\pi]\setminus\{0\}.$ (2)
Proposition 1.
Let
$c_{H}:=\Gamma(2H+1)\sin(\pi H).$
Then, as $\lambda\to 0$,
$f_{H}(\lambda)=c_{H}\,|\lambda|^{1-2H}\bigl(1+o(1)\bigr)$
and
$\partial_{H}f_{H}(\lambda)=f_{H}(\lambda)\Bigl(2\psi(2H+1)+\pi\cot(\pi H)-2\log|\lambda|+o(1)\Bigr),$
where $\psi$ denotes the digamma function. In particular, at $H=3/4$,
$K_{\mathrm{m}}=c_{3/4},\qquad\beta_{\mathrm{m}}=2\psi\!\left(\frac{5}{2}\right)-\pi,$
and $\beta_{\mathrm{m}}$ is the derivative of $\log c_{H}$ at $H=3/4$.
Proof.
Starting from (2), the term $k=0$ dominates as $\lambda\to 0$, while the terms $k\neq 0$ are $O(1)$. Since
$1-\cos\lambda=\frac{\lambda^{2}}{2}+O(\lambda^{4}),$
we obtain
$f_{H}(\lambda)=2c_{H}\left(\frac{\lambda^{2}}{2}+O(\lambda^{4})\right)|\lambda|^{-(2H+1)}\bigl(1+o(1)\bigr)=c_{H}|\lambda|^{1-2H}\bigl(1+o(1)\bigr).$
Differentiating exact series representation (2) with respect to $H$ gives
$\partial_{H}\log f_{H}(\lambda)=\frac{\mathrm{d}}{\mathrm{d}H}\log c_{H}-2\log|\lambda|+o(1)=2\psi(2H+1)+\pi\cot(\pi H)-2\log|\lambda|+o(1),$
which proves the proposition. ∎
At the critical point $H=3/4$ the standard fGn low frequency coefficient is
$K_{\mathrm{m}}:=\Gamma\!\left(\frac{5}{2}\right)\sin\!\left(\frac{3\pi}{4}\right)=\frac{3\sqrt{2\pi}}{8},$ (3)
$\beta_{\mathrm{m}}:=\left.\frac{\mathrm{d}}{\mathrm{d}H}\log\!\bigl(\Gamma(2H+1)\sin(\pi H)\bigr)\right|_{H=3/4}=2\psi\!\left(\frac{5}{2}\right)-\pi=\frac{16}{3}-2\gamma_{\mathrm{E}}-4\log 2-\pi.$ (4)
Thus,
$f_{3/4}(\lambda)=K_{\mathrm{m}}\left|\lambda\right|^{-1/2}\bigl(1+r_{\mathrm{m}}(\lambda)\bigr),\qquad\partial_{H}f_{H}\big|_{H=3/4}(\lambda)=f_{3/4}(\lambda)\bigl(\beta_{\mathrm{m}}-2\log\left|\lambda\right|+\rho_{\mathrm{m}}(\lambda)\bigr),$ (5)
with $r_{\mathrm{m}}(\lambda)\to 0$ and $\rho_{\mathrm{m}}(\lambda)\to 0$ as $\lambda\to 0$.
Fix a critical parameter point
$\vartheta_{\mathrm{m}}:=(\sigma,3/4),\qquad\sigma>0.$
All mfBm score matrices below are evaluated at $\vartheta_{\mathrm{m}}$.
2.2 The mfOU model
Let $X$ be the stationary solution of
$\mathrm{d}X_{t}=-\alpha X_{t}\,\mathrm{d}t+\mathrm{d}M_{t}^{H},\qquad M_{t}^{H}=\sigma B_{t}^{H}+W_{t},\qquad\alpha>0.$
We observe
$X_{n}^{\mathrm{o}}:=(X_{\Delta_{n}},\dots,X_{n\Delta_{n}})^{\top}.$
Let $f_{\Delta}^{\mathrm{o}}(\lambda;\sigma,H,\alpha)$ be the sampled spectral density of the stationary mfOU model (see [3]). At the continuous time frequencies, the stationary mfOU solution has a smooth OU transfer factor multiplying the same fractional singular coefficient $c_{H}=\Gamma(2H+1)\sin(\pi H)$ as in (2). Consequently, the sampled spectral density inherits the same low frequency singular coefficient and logarithmic derivative in the $H$-direction. More precisely, under the normalization used throughout the paper,
$f_{\Delta}^{\mathrm{o}}(\lambda;\sigma,H,\alpha)=c_{H}\,|\lambda|^{1-2H}\bigl(1+o(1)\bigr)\qquad(\lambda\to 0),$
and
$\partial_{H}f_{\Delta}^{\mathrm{o}}(\lambda;\sigma,H,\alpha)=f_{\Delta}^{\mathrm{o}}(\lambda;\sigma,H,\alpha)\Bigl(2\psi(2H+1)+\pi\cot(\pi H)-2\log|\lambda|+o(1)\Bigr).$
Therefore the relevant critical low frequency coefficients are
$K_{\mathrm{o}}:=\Gamma\!\left(\frac{5}{2}\right)\sin\!\left(\frac{3\pi}{4}\right)=\frac{3\sqrt{2\pi}}{8}=K_{\mathrm{m}},$ (6)
and
$\beta_{\mathrm{o}}:=\left.\frac{\mathrm{d}}{\mathrm{d}H}\log\!\bigl(\Gamma(2H+1)\sin(\pi H)\bigr)\right|_{H=3/4}=2\psi\!\left(\frac{5}{2}\right)-\pi=\frac{16}{3}-2\gamma_{\mathrm{E}}-4\log 2-\pi=\beta_{\mathrm{m}}.$ (7)
Only the induced critical score profiles are used below, so no additional second-order expansion of $f_{\Delta}^{\mathrm{o}}$ is required in this paper.
Fix $\vartheta_{\mathrm{o}}:=(\sigma,3/4,\alpha),\sigma>0,\alpha>0$, all mfOU score matrices below are evaluated at $\vartheta_{\mathrm{o}}$.
2.3 Critical covariance constants
Define
$\Xi_{\sigma\sigma}^{\mathrm{m}}$ $:=\frac{4\sigma^{2}K_{\mathrm{m}}^{2}}{\pi}=\frac{9\sigma^{2}}{8},$ $\Xi_{\sigma H}^{\mathrm{m}}$ $:=\frac{2\sigma^{3}K_{\mathrm{m}}^{2}}{\pi}=\frac{9\sigma^{3}}{16},$ $\Xi_{HH}^{\mathrm{m}}$ $:=\frac{4\sigma^{4}K_{\mathrm{m}}^{2}}{3\pi}=\frac{3\sigma^{4}}{8},$ (8)
$\Xi_{\sigma\sigma}^{\mathrm{o}}$ $:=\frac{4\sigma^{2}K_{\mathrm{o}}^{2}}{\pi}=\frac{9\sigma^{2}}{8},$ $\Xi_{\sigma H}^{\mathrm{o}}$ $:=\frac{2\sigma^{3}K_{\mathrm{o}}^{2}}{\pi}=\frac{9\sigma^{3}}{16},$ $\Xi_{HH}^{\mathrm{o}}$ $:=\frac{4\sigma^{4}K_{\mathrm{o}}^{2}}{3\pi}=\frac{3\sigma^{4}}{8},$ $\Xi_{\alpha\alpha}^{\mathrm{o}}$ $:=\frac{1}{\alpha}.$ (9)
Set
$\Gamma_{ab}^{\mathrm{m}}:=\frac{1}{2}\Xi_{ab}^{\mathrm{m}},\qquad\Gamma_{ab}^{\mathrm{o}}:=\frac{1}{2}\Xi_{ab}^{\mathrm{o}}.$
The critical covariance matrices of mfBm and mfOU are
$\Gamma_{\mathrm{m}}^{\mathrm{crit}}:=\begin{pmatrix}\Gamma_{\sigma\sigma}^{\mathrm{m}}&\Gamma_{\sigma H}^{\mathrm{m}}\\ \Gamma_{\sigma H}^{\mathrm{m}}&\Gamma_{HH}^{\mathrm{m}}\end{pmatrix}=\begin{pmatrix}\dfrac{9\sigma^{2}}{16}&\dfrac{9\sigma^{3}}{32}\\[4.30554pt] \dfrac{9\sigma^{3}}{32}&\dfrac{3\sigma^{4}}{16}\end{pmatrix}$, (10)
and
$\Gamma_{\mathrm{o}}^{\mathrm{crit}}:=\begin{pmatrix}\Gamma_{\sigma\sigma}^{\mathrm{o}}&\Gamma_{\sigma H}^{\mathrm{o}}&0\\ \Gamma_{\sigma H}^{\mathrm{o}}&\Gamma_{HH}^{\mathrm{o}}&0\\ 0&0&\Gamma_{\alpha\alpha}^{\mathrm{o}}\end{pmatrix}=\begin{pmatrix}\dfrac{9\sigma^{2}}{16}&\dfrac{9\sigma^{3}}{32}&0\\[4.30554pt] \dfrac{9\sigma^{3}}{32}&\dfrac{3\sigma^{4}}{16}&0\\[4.30554pt] 0&0&\dfrac{1}{2\alpha}\end{pmatrix}$. (11)
In both models,
$\frac{\Gamma_{\sigma H}^{\bullet}}{\sqrt{\Gamma_{\sigma\sigma}^{\bullet}\Gamma_{HH}^{\bullet}}}=\frac{\sqrt{3}}{2}<1,\qquad\bullet\in\{\mathrm{m},\mathrm{o}\}.$ (12)
Hence the critical $(\sigma,H)$-block is already non-degenerate after the first transformation. The low frequency reductions and the explicit one-dimensional integrals leading to (8)–(9) are collected in Appendix A.
Corollary 2.
The matrix $\Gamma_{\mathrm{m}}^{\mathrm{crit}}$ is positive definite, and so is the $2\times 2$ upper-left block of $\Gamma_{\mathrm{o}}^{\mathrm{crit}}$.
Proof.
The correlation identity (12) gives
$\det\Gamma_{\mathrm{m}}^{\mathrm{crit}}>0,\qquad\det\!\left(\Gamma_{\mathrm{o}}^{\mathrm{crit}}\big|_{\{\sigma,H\}}\right)>0.$
Remark 3.
The proposition 1 together with the reductions in Appendix A shows that the critical coefficients come directly from the near zero behavior of the spectral density. $K_{\mathrm{m}}$ and $\beta_{\mathrm{m}}$ are obtained by evaluating the coefficient $c_{H}=\Gamma(2H+1)\sin(\pi H)$ and its logarithmic derivative at $H=3/4$. The corresponding constants $K_{\mathrm{o}}$ and $\beta_{\mathrm{o}}$ are obtained analogously for the sampled mfOU density. With these closed forms, every leading LAN constant (see theorems 9 and 10) is explicit
$K_{\mathrm{m}}=K_{\mathrm{o}}=\frac{3\sqrt{2\pi}}{8},\qquad\beta_{\mathrm{m}}=\beta_{\mathrm{o}}=\frac{16}{3}-2\gamma_{\mathrm{E}}-4\log 2-\pi,$
$\Gamma_{\mathrm{m}}^{\mathrm{crit}}\big|_{\{\sigma,H\}}=\Gamma_{\mathrm{o}}^{\mathrm{crit}}\big|_{\{\sigma,H\}}=\begin{pmatrix}\dfrac{9\sigma^{2}}{16}&\dfrac{9\sigma^{3}}{32}\\[4.30554pt] \dfrac{9\sigma^{3}}{32}&\dfrac{3\sigma^{4}}{16}\end{pmatrix},\qquad I_{H}^{\mathrm{eff,m}}=I_{H}^{\mathrm{eff,o}}=\frac{3\sigma^{4}}{64}.$
where
$I_{H}^{\mathrm{eff},\bullet}=\Gamma_{HH}^{\bullet}-\Gamma_{H\sigma}^{\bullet}\bigl(\Gamma_{\sigma\sigma}^{\bullet}\bigr)^{-1}\Gamma_{\sigma H}^{\bullet},\qquad\bullet\in\{\mathrm{m},\mathrm{o}\}.$
Thus the critical LAN expansion and the score-test normalization do not contain any unnamed leading constants.
2.4 Exact score functions
This subsection establishes the exact score representations at the critical point. These formulas will be used in the next section to derive the trace asymptotics, score CLTs and LAN expansions.
Proposition 4.
Let $\Sigma_{n}(\theta)$ be the relevant covariance matrix and define
$Z_{n}:=\Sigma_{n}(\theta)^{-1/2}X_{n}\sim\mathcal{N}(0,I_{n}).$
Then every Gaussian score function $S_{i,n}(\theta)$ has the exact representation
$S_{i,n}=\frac{1}{2}Q_{n}(M_{i,n}),\qquad M_{i,n}:=\Sigma_{n}(\theta)^{-1/2}\,\partial_{\theta_{i}}\Sigma_{n}(\theta)\,\Sigma_{n}(\theta)^{-1/2}$ (13)
where $Q_{n}(\cdot)$ is defined in equation (1).
Proof.
The Gaussian score identity gives
$S_{i,n}(\theta)=\partial_{\theta_{i}}\ell_{n}(\theta)=-\frac{1}{2}\mathrm{tr}\!\bigl(\Sigma_{n}(\theta)^{-1}\partial_{\theta_{i}}\Sigma_{n}(\theta)\bigr)+\frac{1}{2}X_{n}^{\top}\Sigma_{n}(\theta)^{-1}(\partial_{\theta_{i}}\Sigma_{n}(\theta))\Sigma_{n}(\theta)^{-1}X_{n}.$
Using $\mathbb{E}[X_{n}^{\top}AX_{n}]=\mathrm{tr}(A\Sigma_{n}(\theta))$ yields (13). ∎
We now rewrite the raw score functions at the critical boundary $H=3/4$ into exact decompositions that isolate the singular $\sigma$-direction in the $H$-score. This yields the transformed $H$-scores. For mfBm define
$C_{n}^{\mathrm{m}}:=M_{\sigma,n}^{\mathrm{m}},\qquad M_{H,n}^{\mathrm{m}}:=\Sigma_{n}^{\mathrm{m}}(\theta)^{-1/2}\,\partial_{H}\Sigma_{n}^{\mathrm{m}}(\theta)\,\Sigma_{n}^{\mathrm{m}}(\theta)^{-1/2},$ (14)
all evaluated at $\theta=\vartheta_{\mathrm{m}}$. Since
$\partial_{H}\bigl(\sigma^{2}\Delta_{n}^{2H-1}\bigr)=2\sigma^{2}\Delta_{n}^{2H-1}\log\Delta_{n}=-2\sigma^{2}\Delta_{n}^{2H-1}L_{n},$
one gets the exact splitting
$S_{H,n}^{\mathrm{m}}=-\sigma L_{n}\,S_{\sigma,n}^{\mathrm{m}}+R_{H,n}^{\mathrm{m}},\qquad R_{H,n}^{\mathrm{m}}=\frac{1}{2}Q_{n}(D_{n}^{\mathrm{m}}),\qquad D_{n}^{\mathrm{m}}:=M_{H,n}^{\mathrm{m}}+\sigma L_{n}\,C_{n}^{\mathrm{m}}.$ (15)
The positive sign in the definition of $D_{n}^{\mathrm{m}}$ comes from replacing $\log\Delta_{n}$ by $L_{n}$; this sign convention keeps the critical decomposition unambiguous.
For mfOU define
$C_{n}^{\mathrm{o}}:=M_{\sigma,n}^{\mathrm{o}},\qquad D_{n}^{\mathrm{o}}:=M_{H,n}^{\mathrm{o}}+\sigma L_{n}\,M_{\sigma,n}^{\mathrm{o}},\qquad A_{n}^{\mathrm{o}}:=M_{\alpha,n}^{\mathrm{o}},$
all evaluated at $\theta=\vartheta_{\mathrm{o}}$. Then
$S_{H,n}^{\mathrm{o}}=-\sigma L_{n}\,S_{\sigma,n}^{\mathrm{o}}+R_{H,n}^{\mathrm{o}},\qquad R_{H,n}^{\mathrm{o}}=\frac{1}{2}Q_{n}(D_{n}^{\mathrm{o}}),\qquad S_{\alpha,n}^{\mathrm{o}}=\frac{1}{2}Q_{n}(A_{n}^{\mathrm{o}}).$ (16)
Accordingly, we refer to $(S^{m}_{\sigma,n},R^{m}_{H,n})^{\top}$ in the mfBm case and $(S^{o}_{\sigma,n},R^{o}_{H,n},S^{o}_{\alpha,n})^{\top}$ in the mfOU case as the transformed score vectors.
Here, we define the term critical profile as the leading reduced frequency-domain symbol of the corresponding score at the critical point $H=3/4$.
Critical profiles.
Let
$w_{\mathrm{m}}(u):=\frac{\eta_{\mathrm{m}}\left|u\right|^{-1/2}}{1+\eta_{\mathrm{m}}\left|u\right|^{-1/2}},\qquad\eta_{\mathrm{m}}:=\sigma^{2}K_{\mathrm{m}},$ (17)
and
$w_{\mathrm{o}}(u):=\frac{\eta_{\mathrm{o}}\left|u\right|^{-1/2}}{1+\eta_{\mathrm{o}}\left|u\right|^{-1/2}},\qquad\eta_{\mathrm{o}}:=\sigma^{2}K_{\mathrm{o}}.$ (18)
The critical score profiles are
$g_{\sigma}^{\mathrm{m}}(u):=\frac{2}{\sigma}w_{\mathrm{m}}(u),\qquad g_{\sigma}^{\mathrm{o}}(u):=\frac{2}{\sigma}w_{\mathrm{o}}(u),$ (19)
and
$g_{H,n}^{\mathrm{m}}(u):=w_{\mathrm{m}}(u)\bigl(2L_{n}+\beta_{\mathrm{m}}-2\log\left|u\right|\bigr),\qquad g_{H,n}^{\mathrm{o}}(u):=w_{\mathrm{o}}(u)\bigl(2L_{n}+\beta_{\mathrm{o}}-2\log\left|u\right|\bigr).$ (20)
For the mfOU drift parameter,
$g_{\alpha}^{\mathrm{o}}(u):=-\frac{2\alpha}{\alpha^{2}+u^{2}}.$ (21)
The additional factor $2L_{n}$ in (20) survives after removing the explicit $\sigma L_{n}\,S_{\sigma,n}$ term; it comes from the derivative of the low frequency singularity itself.
Lemma 5.
As $\left|u\right|\to\infty$,
$w_{\mathrm{m}}(u)\sim\eta_{\mathrm{m}}\left|u\right|^{-1/2},\qquad w_{\mathrm{o}}(u)\sim\eta_{\mathrm{o}}\left|u\right|^{-1/2}.$
Consequently,
$\bigl(g_{\sigma}^{\mathrm{m}}(u)\bigr)^{2}$ $\sim\frac{4\eta_{\mathrm{m}}^{2}}{\sigma^{2}}\left|u\right|^{-1},$ $\bigl(g_{\sigma}^{\mathrm{o}}(u)\bigr)^{2}$ $\sim\frac{4\eta_{\mathrm{o}}^{2}}{\sigma^{2}}\left|u\right|^{-1},$
$\bigl(g_{H,n}^{\mathrm{m}}(u)\bigr)^{2}$ $\sim\eta_{\mathrm{m}}^{2}\left|u\right|^{-1}\bigl(2L_{n}+\beta_{\mathrm{m}}-2\log\left|u\right|\bigr)^{2},$ $\bigl(g_{H,n}^{\mathrm{o}}(u)\bigr)^{2}$ $\sim\eta_{\mathrm{o}}^{2}\left|u\right|^{-1}\bigl(2L_{n}+\beta_{\mathrm{o}}-2\log\left|u\right|\bigr)^{2},$
and
$g_{\sigma}^{\mathrm{m}}(u)g_{H,n}^{\mathrm{m}}(u)$ $\sim\frac{2\eta_{\mathrm{m}}^{2}}{\sigma}\left|u\right|^{-1}\bigl(2L_{n}+\beta_{\mathrm{m}}-2\log\left|u\right|\bigr),$
$g_{\sigma}^{\mathrm{o}}(u)g_{H,n}^{\mathrm{o}}(u)$ $\sim\frac{2\eta_{\mathrm{o}}^{2}}{\sigma}\left|u\right|^{-1}\bigl(2L_{n}+\beta_{\mathrm{o}}-2\log\left|u\right|\bigr),$
$g_{\alpha}^{\mathrm{o}}(u)$ $\sim-2\alpha\left|u\right|^{-2}.$
Proof.
This is immediate from (17)–(21). ∎
3 Main Results
We now state the central limit theorems for the transformed score vectors in equations (15) and (16) at the critical boundary $H=3/4$. Throughout this subsection, let
$L_{n}:=\log(1/\Delta_{n}),\qquad\Delta_{n}\downarrow 0,\qquad n\Delta_{n}\to\infty,$
and all score matrices and score functions are evaluated at the critical parameter point. For the mfBm model we write
$\vartheta_{m}:=(\sigma,3/4),\qquad\sigma>0,$
and for the mfOU model,
$\vartheta_{o}:=(\sigma,3/4,\alpha),\qquad\sigma>0,\ \alpha>0.$
Recall that the $H$-score admits the exact decomposition
$S^{m}_{H,n}=-\sigma L_{n}S^{m}_{\sigma,n}+R^{m}_{H,n},\qquad S^{o}_{H,n}=-\sigma L_{n}S^{o}_{\sigma,n}+R^{o}_{H,n},$
where $R^{m}_{H,n}$ and $R^{o}_{H,n}$ denote the transformed $H$-scores after removing the explicit singular linear term in the $\sigma$-direction.
The next two theorems show that, after normalization by the exact critical rates, the transformed score vectors are asymptotically Gaussian.
Theorem 6.
Consider the mixed fractional Brownian motion model at the critical point $\vartheta_{m}=(\sigma,3/4)$. Then, as $n\to\infty$,
$\left(\frac{S_{\sigma,n}^{\mathrm{m}}}{\sqrt{n\Delta_{n}L_{n}}},\frac{R_{H,n}^{\mathrm{m}}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}\right)\xrightarrow[]{d}\mathcal{N}\bigl(0,\Gamma_{\mathrm{m}}^{\mathrm{crit}}\bigr)$
where $\Gamma_{\mathrm{m}}^{\mathrm{crit}}$ is defined by equation (10). In particular, the critical fluctuation scale of the $\sigma$-score is $(n\Delta_{n}L_{n})^{1/2}$, whereas the critical fluctuation scale of the transformed $H$-score is $(n\Delta_{n}L_{n}^{3})^{1/2}$.
Theorem 7.
Consider the mixed fractional Ornstein–Uhlenbeck model at the critical point $\vartheta_{o}=(\sigma,3/4,\alpha)$. Then, as $n\to\infty$,
$\left(\frac{S_{\sigma,n}^{\mathrm{o}}}{\sqrt{n\Delta_{n}L_{n}}},\frac{R_{H,n}^{\mathrm{o}}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}},\frac{S_{\alpha,n}^{\mathrm{o}}}{\sqrt{n\Delta_{n}}}\right)\xrightarrow[]{d}\mathcal{N}\bigl(0,\Gamma_{\mathrm{o}}^{\mathrm{crit}}\bigr).$
where $\Gamma_{\mathrm{o}}^{\mathrm{crit}}$ is defined by equation (11). Hence, at the critical point, the $(\sigma,H)$-block and the $\alpha$-direction are asymptotically orthogonal at the leading order.
The previous theorems are formulated in terms of the transformed $H$-score $R_{H,n}$ rather than the raw $H$-score $S_{H,n}$. By (15) and (16), the two score functions are related by the following triangular changes of coordinates.
Proposition 8.
Let
$L_{n}^{\mathrm{m}}:=\begin{pmatrix}1&0\\ \sigma L_{n}&1\end{pmatrix},\qquad L_{n}^{\mathrm{o}}:=\begin{pmatrix}1&0&0\\ \sigma L_{n}&1&0\\ 0&0&1\end{pmatrix}$.
Then
$L_{n}^{\mathrm{m}}\begin{pmatrix}S_{\sigma,n}^{\mathrm{m}}\\ S_{H,n}^{\mathrm{m}}\end{pmatrix}=\begin{pmatrix}S_{\sigma,n}^{\mathrm{m}}\\ R_{H,n}^{\mathrm{m}}\end{pmatrix},\qquad L_{n}^{\mathrm{o}}\begin{pmatrix}S_{\sigma,n}^{\mathrm{o}}\\ S_{H,n}^{\mathrm{o}}\\ S_{\alpha,n}^{\mathrm{o}}\end{pmatrix}=\begin{pmatrix}S_{\sigma,n}^{\mathrm{o}}\\ R_{H,n}^{\mathrm{o}}\\ S_{\alpha,n}^{\mathrm{o}}\end{pmatrix}$.
Thus, at the critical boundary $H=3/4$, the explicit singular coupling in the $\sigma$-direction is removed by a single triangular transformation, and no second projection is required.
Proof.
This is a direct restatement of (15) and (16) in matrix form. Moreover, by (12),
$\frac{\Gamma^{\bullet}_{\sigma H}}{\sqrt{\Gamma^{\bullet}_{\sigma\sigma}\Gamma^{\bullet}_{HH}}}=\frac{\sqrt{3}}{2}<1,\qquad\bullet\in\{\mathrm{m},\mathrm{o}\},$
so the critical $(\sigma,H)$-block is already non-degenerate after the first transformation. Therefore no second projection is required. ∎
For later use in the LAN expansion, we introduce the corresponding critical rate matrices
$r^{-1}_{n,m}:=\begin{pmatrix}(n\Delta_{n}L_{n})^{-1/2}&\sigma(n\Delta_{n}L_{n})^{-1/2}\\ 0&(n\Delta_{n})^{-1/2}L_{n}^{-3/2}\end{pmatrix}$,
and
$r^{-1}_{n,o}:=\begin{pmatrix}(n\Delta_{n}L_{n})^{-1/2}&\sigma(n\Delta_{n}L_{n})^{-1/2}&0\\ 0&(n\Delta_{n})^{-1/2}L_{n}^{-3/2}&0\\ 0&0&(n\Delta_{n})^{-1/2}\end{pmatrix}$.
These matrices encode exactly the critical local scales of the parameters and the first-order coupling between the $\sigma$- and $H$-directions.
Theorem 9.
Let $\ell_{n}^{\mathrm{m}}$ be the Gaussian log-likelihood of the mfBm increment experiment and define
$\Xi_{n}^{\mathrm{m}}:=\begin{pmatrix}\dfrac{S_{\sigma,n}^{\mathrm{m}}}{\sqrt{n\Delta_{n}L_{n}}}\\[3.44444pt] \dfrac{R_{H,n}^{\mathrm{m}}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}\end{pmatrix},\qquad\vartheta_{\mathrm{m},n}(h):=\vartheta_{\mathrm{m}}+r_{n,\mathrm{m}}^{-1}h,\qquad h\in\mathbb{R}^{2}.$
Then for every fixed $h\in\mathbb{R}^{2}$,
$\ell_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m},n}(h))-\ell_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})=h^{\top}\Xi_{n}^{\mathrm{m}}-\frac{1}{2}h^{\top}\Gamma_{\mathrm{m}}^{\mathrm{crit}}h+o_{P_{\vartheta_{\mathrm{m}}}}(1).$
In particular, the mfBm experiment is LAN at $\vartheta_{\mathrm{m}}$ with asymptotic information $\Gamma_{\mathrm{m}}^{\mathrm{crit}}$.
Theorem 10.
Let $\ell_{n}^{\mathrm{o}}$ be the Gaussian log-likelihood of the stationary mfOU experiment and define
$\Xi_{n}^{\mathrm{o}}:=\begin{pmatrix}\dfrac{S_{\sigma,n}^{\mathrm{o}}}{\sqrt{n\Delta_{n}L_{n}}}\\[3.44444pt] \dfrac{R_{H,n}^{\mathrm{o}}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}\\[3.44444pt] \dfrac{S_{\alpha,n}^{\mathrm{o}}}{\sqrt{n\Delta_{n}}}\end{pmatrix},\qquad\vartheta_{\mathrm{o},n}(h):=\vartheta_{\mathrm{o}}+r_{n,\mathrm{o}}^{-1}h,\qquad h\in\mathbb{R}^{3}.$
Then for every fixed $h\in\mathbb{R}^{3}$,
$\ell_{n}^{\mathrm{o}}(\vartheta_{\mathrm{o},n}(h))-\ell_{n}^{\mathrm{o}}(\vartheta_{\mathrm{o}})=h^{\top}\Xi_{n}^{\mathrm{o}}-\frac{1}{2}h^{\top}\Gamma_{\mathrm{o}}^{\mathrm{crit}}h+o_{P_{\vartheta_{\mathrm{o}}}}(1).$
In particular, the mfOU experiment is LAN at $\vartheta_{\mathrm{o}}$ with asymptotic information $\Gamma_{\mathrm{o}}^{\mathrm{crit}}$.
4 Proof of the Score CLTs
This section proves the score central limit theorems. We first establish the critical trace asymptotics for the score matrices and then combine them with the Gaussian quadratic-form CLT (see lemma 13) to derive the limiting normal laws.
4.1 Critical trace asymptotics
Proposition 11.
Consider the mixed fractional Brownian motion model at the critical point
$\vartheta_{\mathrm{m}}=(\sigma,3/4),\qquad\sigma>0,$
under the high-frequency asymptotic regime
$\Delta_{n}\downarrow 0,\qquad n\Delta_{n}\to\infty,\qquad L_{n}:=\log(1/\Delta_{n}).$
Then the following trace asymptotics hold:
$\mathrm{tr}\bigl((C_{n}^{\mathrm{m}})^{2}\bigr)$ $\sim\Xi_{\sigma\sigma}^{\mathrm{m}}\,n\Delta_{n}L_{n},$ (22)
$\mathrm{tr}(C_{n}^{\mathrm{m}}D_{n}^{\mathrm{m}})$ $\sim\Xi_{\sigma H}^{\mathrm{m}}\,n\Delta_{n}L_{n}^{2},$ (23)
$\mathrm{tr}\bigl((D_{n}^{\mathrm{m}})^{2}\bigr)$ $\sim\Xi_{HH}^{\mathrm{m}}\,n\Delta_{n}L_{n}^{3},$ (24)
and
$\left\lVert C_{n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}=O(1),\qquad\left\lVert D_{n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}=O(L_{n}).$ (25)
Here and throughout, the notation is as in Section 2.
Proof.
At $H=3/4$, after the low frequency rescaling $\lambda=\Delta_{n}u$, the critical profiles corresponding to $C_{n}^{\mathrm{m}}$ and $D_{n}^{\mathrm{m}}$ are $g_{\sigma}^{\mathrm{m}}(u)$ and $g_{H,n}^{\mathrm{m}}(u)$. Therefore,
$\mathrm{tr}\bigl((C_{n}^{\mathrm{m}})^{2}\bigr),\qquad\mathrm{tr}(C_{n}^{\mathrm{m}}D_{n}^{\mathrm{m}}),\qquad\mathrm{tr}\bigl((D_{n}^{\mathrm{m}})^{2}\bigr)$
are reduced to the corresponding one-dimensional critical integrals via the standard Toeplitz trace reduction (see [2]).
For (22), split the integral into $\left|u\right|\leq 1$ and $1\leq\left|u\right|\leq 1/\Delta_{n}$. The bounded part is $O(1)$. By lemma 5,
$\int_{1}^{1/\Delta_{n}}\bigl(g_{\sigma}^{\mathrm{m}}(u)\bigr)^{2}\,\mathrm{d}u\sim\frac{4\eta_{\mathrm{m}}^{2}}{\sigma^{2}}\int_{1}^{1/\Delta_{n}}u^{-1}\,\mathrm{d}u=\frac{4\eta_{\mathrm{m}}^{2}}{\sigma^{2}}L_{n}.$
Doubling by evenness and multiplying by $n\Delta_{n}/(2\pi)$ gives
$\int_{-1/\Delta_{n}}^{1/\Delta_{n}}\bigl(g_{\sigma}^{\mathrm{m}}(u)\bigr)^{2}\,\mathrm{d}u\sim\frac{8\eta_{\mathrm{m}}^{2}}{\sigma^{2}}L_{n},$
hence (22).
For (23), the bounded region contributes only $O(L_{n})$ and is therefore negligible after division by $L_{n}^{2}$. On $1,1/\Delta_{n}]$,
$\int_{1}^{1/\Delta_{n}}g_{\sigma}^{\mathrm{m}}(u)g_{H,n}^{\mathrm{m}}(u)\,\mathrm{d}u\sim\frac{2\eta_{\mathrm{m}}^{2}}{\sigma}\int_{1}^{1/\Delta_{n}}u^{-1}\bigl(2L_{n}+\beta_{\mathrm{m}}-2\log u\bigr)\,\mathrm{d}u.$
With $t=\log u$,
$\int_{1}^{1/\Delta_{n}}u^{-1}\bigl(2L_{n}+\beta_{\mathrm{m}}-2\log u\bigr)\,\mathrm{d}u=\int_{0}^{L_{n}}(2L_{n}+\beta_{\mathrm{m}}-2t)\,\mathrm{d}t=L_{n}^{2}+\beta_{\mathrm{m}}L_{n}.$
After doubling by evenness, (23) follows.
For (24),
$\int_{1}^{1/\Delta_{n}}\bigl(g_{H,n}^{\mathrm{m}}(u)\bigr)^{2}\,\mathrm{d}u\sim\eta_{\mathrm{m}}^{2}\int_{1}^{1/\Delta_{n}}u^{-1}\bigl(2L_{n}+\beta_{\mathrm{m}}-2\log u\bigr)^{2}\,\mathrm{d}u.$
Again $t=\log u$ gives
$\int_{1}^{1/\Delta_{n}}u^{-1}\bigl(2L_{n}+\beta_{\mathrm{m}}-2\log u\bigr)^{2}\,\mathrm{d}u$ $=\int_{0}^{L_{n}}(2L_{n}+\beta_{\mathrm{m}}-2t)^{2}\,\mathrm{d}t$
$=\frac{4}{3}L_{n}^{3}+2\beta_{\mathrm{m}}L_{n}^{2}+\beta_{\mathrm{m}}^{2}L_{n}.$
Hence the leading even integral has coefficient $8\eta_{\mathrm{m}}^{2}/3$, which yields (24).
For $C_{n}^{m}$, the operator bound follows from the boundedness of the corresponding reduced symbol. For $D_{n}^{m}$, however, the reduced profile $g_{H,n}$ is not bounded near the origin because of the logarithmic term. Hence the bound $D_{n}^{m}||_{op}=O(L_{n})$ requires a separate argument at the level of the original sandwiched symbol, which we omit here and prove below. ∎
Proposition 12.
Consider the mixed fractional Ornstein–Uhlenbeck model at the critical point
$\vartheta_{\mathrm{o}}=(\sigma,3/4,\alpha),\qquad\sigma>0,\ \alpha>0,$
under the high-frequency asymptotic regime
$\Delta_{n}\downarrow 0,\qquad n\Delta_{n}\to\infty,\qquad L_{n}:=\log(1/\Delta_{n}).$
Then the following trace asymptotics hold:
$\mathrm{tr}\bigl((C_{n}^{\mathrm{o}})^{2}\bigr)$ $\sim\Xi_{\sigma\sigma}^{\mathrm{o}}\,n\Delta_{n}L_{n},$ (26)
$\mathrm{tr}(C_{n}^{\mathrm{o}}D_{n}^{\mathrm{o}})$ $\sim\Xi_{\sigma H}^{\mathrm{o}}\,n\Delta_{n}L_{n}^{2},$ (27)
$\mathrm{tr}\bigl((D_{n}^{\mathrm{o}})^{2}\bigr)$ $\sim\Xi_{HH}^{\mathrm{o}}\,n\Delta_{n}L_{n}^{3},$ (28)
$\mathrm{tr}\bigl((A_{n}^{\mathrm{o}})^{2}\bigr)$ $\sim\Xi_{\alpha\alpha}^{\mathrm{o}}\,n\Delta_{n},$ (29)
and
$\left\lVert C_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O(1),\qquad\left\lVert D_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O(L_{n}),\qquad\left\lVert A_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O(1).$ (30)
Moreover,
$\mathrm{tr}(C_{n}^{\mathrm{o}}A_{n}^{\mathrm{o}})=O(n\Delta_{n}),\qquad\mathrm{tr}(D_{n}^{\mathrm{o}}A_{n}^{\mathrm{o}})=O(n\Delta_{n}L_{n}).$ (31)
Here and throughout, the notation is as in Section 2.
Proof.
For the $(\sigma,H)$-block, the critical rescaling of the sampled mfOU spectral density yields the same integrals as in the mfBm case (see [3]), with $\eta_{\mathrm{m}}$ replaced by $\eta_{\mathrm{o}}$ (see Appendix A for the explicit low frequency reduction). This proves (26)–(28).
For the drift score,
$\int_{\mathbb{R}}\bigl(g_{\alpha}^{\mathrm{o}}(u)\bigr)^{2}\,\mathrm{d}u=4\alpha^{2}\int_{\mathbb{R}}\frac{\mathrm{d}u}{(\alpha^{2}+u^{2})^{2}}=\frac{2\pi}{\alpha},$
which gives (29) after multiplication by $n\Delta_{n}/(2\pi)$.
For the mixed drift traces, $g_{\sigma}^{\mathrm{o}}$ is bounded near the origin and behaves like $\left|u\right|^{-1/2}$ at infinity, whereas $g_{\alpha}^{\mathrm{o}}(u)$ behaves like $\left|u\right|^{-2}$. Hence
$\int_{\mathbb{R}}\left|g_{\sigma}^{\mathrm{o}}(u)g_{\alpha}^{\mathrm{o}}(u)\right|\,\mathrm{d}u<\infty,$
so $\mathrm{tr}(C_{n}^{\mathrm{o}}A_{n}^{\mathrm{o}})=O(n\Delta_{n})$. Likewise,
$\int_{\mathbb{R}}\left|g_{H,n}^{\mathrm{o}}(u)g_{\alpha}^{\mathrm{o}}(u)\right|\,\mathrm{d}u\leq C(1+L_{n})\int_{\mathbb{R}}\frac{\mathrm{d}u}{1+u^{2}}<\infty,$
which yields $\mathrm{tr}(D_{n}^{\mathrm{o}}A_{n}^{\mathrm{o}})=O(n\Delta_{n}L_{n})$. The operator bounds in (30) follow from the same profile bounds as in the mfBm case. ∎
4.2 CLT for the transformed score vector
The proof is based on the operator/Frobenius criterion for Gaussian quadratic forms.
Lemma 13.
Let $A_{n}$ be real symmetric matrices and let $Z_{n}\sim\mathcal{N}(0,I_{n})$. If
$\mathrm{tr}(A_{n}^{2})\to\tau^{2}\in(0,\infty),\qquad\frac{\left\lVert A_{n}\right\rVert_{\mathrm{op}}}{\left\lVert A_{n}\right\rVert_{\mathrm{F}}}\to 0,$
then
$\frac{1}{2}Q_{n}(A_{n})\xrightarrow[]{d}\mathcal{N}\!\left(0,\frac{\tau^{2}}{2}\right).$
Proof.
See [2]. ∎
Proof of theorem 6.
Fix $u=(u_{1},u_{2})\in\mathbb{R}^{2}$ and define
$A_{n}^{\mathrm{m}}(u):=\frac{u_{1}}{\sqrt{n\Delta_{n}L_{n}}}C_{n}^{\mathrm{m}}+\frac{u_{2}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}D_{n}^{\mathrm{m}}$
where $C_{n}^{m}$ and $D_{n}^{m}$ are defined in (14) and (15). Then
$\frac{u_{1}S_{\sigma,n}^{\mathrm{m}}}{\sqrt{n\Delta_{n}L_{n}}}+\frac{u_{2}R_{H,n}^{\mathrm{m}}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}=\frac{1}{2}Q_{n}\bigl(A_{n}^{\mathrm{m}}(u)\bigr).$
By proposition 11,
$\mathrm{tr}\Bigl(\bigl(A_{n}^{\mathrm{m}}(u)\bigr)^{2}\Bigr)\to u^{\top}\bigl(2\Gamma_{\mathrm{m}}^{\mathrm{crit}}\bigr)u.$
Moreover,
$\left\lVert A_{n}^{\mathrm{m}}(u)\right\rVert_{\mathrm{op}}\leq\frac{|u_{1}|}{\sqrt{n\Delta_{n}L_{n}}}\left\lVert C_{n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}+\frac{|u_{2}|}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}\left\lVert D_{n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}=O\!\left(\frac{1}{\sqrt{n\Delta_{n}L_{n}}}\right),$
whereas $\left\lVert A_{n}^{\mathrm{m}}(u)\right\rVert_{\mathrm{F}}^{2}\to 2u^{\top}\Gamma_{\mathrm{m}}^{\mathrm{crit}}u$. Since $\Gamma_{\mathrm{m}}^{\mathrm{crit}}$ is positive definite, the operator/Frobenius ratio tends to 0 for every $u\neq 0$. By lemma 13,
$\frac{1}{2}Q_{n}\bigl(A_{n}^{\mathrm{m}}(u)\bigr)\xrightarrow[]{d}\mathcal{N}\bigl(0,u^{\top}\Gamma_{\mathrm{m}}^{\mathrm{crit}}u\bigr).$
The Cramér–Wold device proves the theorem. ∎
Proof of theorem 7.
Fix $u=(u_{1},u_{2},u_{3})\in\mathbb{R}^{3}$ and define
$A_{n}^{\mathrm{o}}(u):=\frac{u_{1}}{\sqrt{n\Delta_{n}L_{n}}}C_{n}^{\mathrm{o}}+\frac{u_{2}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}D_{n}^{\mathrm{o}}+\frac{u_{3}}{\sqrt{n\Delta_{n}}}A_{n}^{\mathrm{o}}.$
Then
$\frac{u_{1}S_{\sigma,n}^{\mathrm{o}}}{\sqrt{n\Delta_{n}L_{n}}}+\frac{u_{2}R_{H,n}^{\mathrm{o}}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}+\frac{u_{3}S_{\alpha,n}^{\mathrm{o}}}{\sqrt{n\Delta_{n}}}=\frac{1}{2}Q_{n}\bigl(A_{n}^{\mathrm{o}}(u)\bigr).$
By proposition 12, the diagonal terms contribute $2u^{\top}\Gamma_{\mathrm{o}}^{\mathrm{crit}}u$ at the trace level. The mixed traces with $A_{n}^{\mathrm{o}}$ vanish after normalization because
$\frac{n\Delta_{n}}{\sqrt{n\Delta_{n}L_{n}}\sqrt{n\Delta_{n}}}=L_{n}^{-1/2}\to 0,\qquad\frac{n\Delta_{n}L_{n}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}\sqrt{n\Delta_{n}}}=L_{n}^{-1/2}\to 0.$
Hence
$\mathrm{tr}\Bigl(\bigl(A_{n}^{\mathrm{o}}(u)\bigr)^{2}\Bigr)\to 2u^{\top}\Gamma_{\mathrm{o}}^{\mathrm{crit}}u.$
Also,
$\left\lVert A_{n}^{\mathrm{o}}(u)\right\rVert_{\mathrm{op}}\leq\frac{|u_{1}|}{\sqrt{n\Delta_{n}L_{n}}}\left\lVert C_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}+\frac{|u_{2}|}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}\left\lVert D_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}+\frac{|u_{3}|}{\sqrt{n\Delta_{n}}}\left\lVert A_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O\!\left(\frac{1}{\sqrt{n\Delta_{n}}}\right).$
Since $\Gamma_{\mathrm{o}}^{\mathrm{crit}}$ is positive definite, the Frobenius norm of $A_{n}^{\mathrm{o}}(u)$ stays away from 0 for every $u\neq 0$, and therefore
$\frac{\left\lVert A_{n}^{\mathrm{o}}(u)\right\rVert_{\mathrm{op}}}{\left\lVert A_{n}^{\mathrm{o}}(u)\right\rVert_{\mathrm{F}}}\to 0.$
Lemma 13 and the Cramér–Wold device conclude the proof. ∎
5 Proof of the LAN theorems
In this section we combine the CLTs for the transformed score vectors with a second-order expansion of the Gaussian likelihood ratio. We first state the matrix Taylor bounds. We then linearize the local covariance perturbation. Finally, we identify the linear score term, the quadratic information term and the remainder terms. At the critical point $H=3/4$, a single triangular rate matrix already yields a non-degenerate quadratic term.
5.1 Matrix Taylor expansions
Subsection 5.1 collects the matrix expansions used in the LAN proof.
Lemma 14.
Let $S$ be a real symmetric matrix with $\left\lVert S\right\rVert_{\mathrm{op}}\leq 1/2$. Then
$\log\det(I+S)=\mathrm{tr}(S)-\frac{1}{2}\mathrm{tr}(S^{2})+R_{\log}(S),\qquad\left|R_{\log}(S)\right|\leq C\left\lVert S\right\rVert_{\mathrm{op}}\mathrm{tr}(S^{2}),$
and
$(I+S)^{-1}=I-S+S^{2}+R_{\mathrm{inv}}(S),\qquad R_{\mathrm{inv}}(S)=\sum_{k\geq 3}(-1)^{k}S^{k},$
with
$\left\lVert R_{\mathrm{inv}}(S)\right\rVert_{\mathrm{op}}\leq C\left\lVert S\right\rVert_{\mathrm{op}}^{3},\qquad\mathrm{tr}\bigl(R_{\mathrm{inv}}(S)^{2}\bigr)\leq C\mathrm{tr}(S^{6}).$
Here $C>0$ is a universal constant.
Proof.
Diagonalize $S$. Since every eigenvalue belongs to [1/2,1/2][-1/2,1/2], the scalar Taylor expansions of $\log(1+x)$ and $(1+x)^{-1}$ can be summed over the spectrum term by term (see [3]). ∎
5.2 Critical local perturbation
This subsection linearizes the local covariance perturbation under the critical local parameterization at $H=3/4$. For the local parameter shift $\vartheta_{n}(h)$, we write the sandwiched perturbation $S_{n}(h)$ as its first-order part $B_{n}(h)$ plus a negligible remainder. This is the key for the likelihood expansion in the next subsection.
Proposition 15.
Fix $h\in\mathbb{R}^{2}$ and set
$\vartheta_{\mathrm{m},n}(h)=\vartheta_{\mathrm{m}}+r_{n,\mathrm{m}}^{-1}h,\qquad B_{n}^{\mathrm{m}}(h):=\frac{h_{1}}{\sqrt{n\Delta_{n}L_{n}}}C_{n}^{\mathrm{m}}+\frac{h_{2}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}D_{n}^{\mathrm{m}}.$
Define the sandwiched covariance perturbation
$\mathcal{S}_{n}^{\mathrm{m}}(h):=\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})^{-1/2}\bigl(\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m},n}(h))-\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})\bigr)\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})^{-1/2}.$
Then
$\mathcal{S}_{n}^{\mathrm{m}}(h)=B_{n}^{\mathrm{m}}(h)+\mathcal{R}_{n}^{\mathrm{m}}(h),\qquad\left\lVert\mathcal{S}_{n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}\to 0,\qquad\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)\to 0.$
Fix $h\in\mathbb{R}^{3}$ and set
$\vartheta_{\mathrm{o},n}(h)=\vartheta_{\mathrm{o}}+r_{n,\mathrm{o}}^{-1}h,\qquad B_{n}^{\mathrm{o}}(h):=\frac{h_{1}}{\sqrt{n\Delta_{n}L_{n}}}C_{n}^{\mathrm{o}}+\frac{h_{2}}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}D_{n}^{\mathrm{o}}+\frac{h_{3}}{\sqrt{n\Delta_{n}}}A_{n}^{\mathrm{o}}.$
Define $\mathcal{S}_{n}^{\mathrm{o}}(h)$ analogously. Then
$\mathcal{S}_{n}^{\mathrm{o}}(h)=B_{n}^{\mathrm{o}}(h)+\mathcal{R}_{n}^{\mathrm{o}}(h),\qquad\left\lVert\mathcal{S}_{n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}\to 0,\qquad\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)\to 0.$
Proof.
We start with mfBm. Write $\delta_{n}^{\mathrm{m}}(h)=r_{n,\mathrm{m}}^{-1}h$, so that
$\delta_{\sigma,n}^{\mathrm{m}}=O\bigl((n\Delta_{n}L_{n})^{-1/2}\bigr),\qquad\delta_{H,n}^{\mathrm{m}}=O\bigl((n\Delta_{n})^{-1/2}L_{n}^{-3/2}\bigr).$
A second-order Taylor expansion around $\vartheta_{\mathrm{m}}$ yields
$\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m},n}(h))-\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})=\delta_{\sigma,n}^{\mathrm{m}}\,\partial_{\sigma}\Sigma_{n}^{\mathrm{m}}+\delta_{H,n}^{\mathrm{m}}\,\partial_{H}\Sigma_{n}^{\mathrm{m}}+\frac{1}{2}\sum_{i,j\in\{\sigma,H\}}\delta_{i,n}^{\mathrm{m}}\delta_{j,n}^{\mathrm{m}}\partial_{ij}^{2}\Sigma_{n}^{\mathrm{m}}(\widetilde{\vartheta}_{n}),$
for some $\widetilde{\vartheta}_{n}$ on the segment joining $\vartheta_{\mathrm{m}}$ and $\vartheta_{\mathrm{m},n}(h)$. Sandwiching by $\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})^{-1/2}$ and using the exact decomposition (15) gives
$\mathcal{S}_{n}^{\mathrm{m}}(h)=B_{n}^{\mathrm{m}}(h)+\mathcal{R}_{n}^{\mathrm{m}}(h),\qquad\mathcal{R}_{n}^{\mathrm{m}}(h)=\frac{1}{2}\sum_{i,j\in\{\sigma,H\}}\delta_{i,n}^{\mathrm{m}}\delta_{j,n}^{\mathrm{m}}M_{ij,n}^{\mathrm{m}}(\widetilde{\vartheta}_{n}),$
where
$M_{ij,n}^{\mathrm{m}}(\theta):=\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})^{-1/2}\partial_{ij}^{2}\Sigma_{n}^{\mathrm{m}}(\theta)\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})^{-1/2}.$
Differentiating
$\Sigma_{n}^{\mathrm{m}}(\sigma,H)=\Delta_{n}\Bigl(I_{n}+\sigma^{2}\Delta_{n}^{2H-1}\mathcal{T}_{n}(f_{H})\Bigr)$
term by term, and using the higher-order derivative bounds proved later in proposition 21, we obtain uniformly on shrinking neighborhoods of $\vartheta_{\mathrm{m}}$
$\left\lVert M_{\sigma\sigma,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O(1),$ $\left\lVert M_{\sigma H,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O(L_{n}),$ $\left\lVert M_{HH,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O(L_{n}^{2}),$
$\mathrm{tr}\bigl((M_{\sigma\sigma,n}^{\mathrm{m}})^{2}\bigr)$ $=O(n\Delta_{n}L_{n}),$ $\mathrm{tr}\bigl((M_{\sigma H,n}^{\mathrm{m}})^{2}\bigr)$ $=O(n\Delta_{n}L_{n}^{3}),$ $\mathrm{tr}\bigl((M_{HH,n}^{\mathrm{m}})^{2}\bigr)$ $=O(n\Delta_{n}L_{n}^{5}).$
Equivalently, each second-order Taylor contribution is already below the LAN scale:
$(\delta_{\sigma,n}^{\mathrm{m}})^{2}\left\lVert M_{\sigma\sigma,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$
$|\delta_{\sigma,n}^{\mathrm{m}}\delta_{H,n}^{\mathrm{m}}|\left\lVert M_{\sigma H,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$
$(\delta_{H,n}^{\mathrm{m}})^{2}\left\lVert M_{HH,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$
and, at the Frobenius level,
$(\delta_{\sigma,n}^{\mathrm{m}})^{4}\mathrm{tr}\bigl((M_{\sigma\sigma,n}^{\mathrm{m}})^{2}\bigr)$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$
$(\delta_{\sigma,n}^{\mathrm{m}}\delta_{H,n}^{\mathrm{m}})^{2}\mathrm{tr}\bigl((M_{\sigma H,n}^{\mathrm{m}})^{2}\bigr)$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$
$(\delta_{H,n}^{\mathrm{m}})^{4}\mathrm{tr}\bigl((M_{HH,n}^{\mathrm{m}})^{2}\bigr)$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr).$
The $HH$-term is generated by the three critical profiles
$\Delta_{n}^{3/2}L_{n}^{2}\mathcal{T}_{n}(f_{3/4}),\qquad\Delta_{n}^{3/2}L_{n}\mathcal{T}_{n}(\partial_{H}f_{H}|_{H=3/4}),\qquad\Delta_{n}^{3/2}\mathcal{T}_{n}(\partial_{HH}^{2}f_{H}|_{H=3/4}),$
and the last one has low frequency order $|\lambda|^{-1/2}(1+|\log|\lambda||^{2})$, so it falls under exactly the same critical trace mechanism as the first-order score matrices. Therefore
$\left\lVert\mathcal{R}_{n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),\qquad\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr)\to 0.$
Since equation (25) implies $\left\lVert B_{n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}\to 0$, we also obtain $\left\lVert\mathcal{S}_{n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}\to 0$.
For mfOU we write $\delta_{n}^{\mathrm{o}}(h)=r_{n,\mathrm{o}}^{-1}h$ and repeat the same argument. The local Taylor expansion gives
$\mathcal{S}_{n}^{\mathrm{o}}(h)=B_{n}^{\mathrm{o}}(h)+\mathcal{R}_{n}^{\mathrm{o}}(h),\qquad\mathcal{R}_{n}^{\mathrm{o}}(h)=\frac{1}{2}\sum_{i,j\in\{\sigma,H,\alpha\}}\delta_{i,n}^{\mathrm{o}}\delta_{j,n}^{\mathrm{o}}M_{ij,n}^{\mathrm{o}}(\widetilde{\vartheta}_{n}).$
Because the sampled mfOU symbol has the same critical $|u|^{-1/2}$ singularity as in the mfBm case, each $H$-derivative contributes one logarithmic factor, whereas $\alpha$-derivatives only act on the smooth OU correction and therefore preserve the integrability class. Thus
$\left\lVert M_{\sigma\sigma,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(1),$ $\left\lVert M_{\sigma H,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(L_{n}),$ $\left\lVert M_{HH,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(L_{n}^{2}),$
$\left\lVert M_{\sigma\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(1),$ $\left\lVert M_{H\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(L_{n}),$ $\left\lVert M_{\alpha\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(1),$
while
$\mathrm{tr}\bigl((M_{\sigma\sigma,n}^{\mathrm{o}})^{2}\bigr)$ $=O(n\Delta_{n}L_{n}),$ $\mathrm{tr}\bigl((M_{\sigma H,n}^{\mathrm{o}})^{2}\bigr)$ $=O(n\Delta_{n}L_{n}^{3}),$ $\mathrm{tr}\bigl((M_{HH,n}^{\mathrm{o}})^{2}\bigr)$ $=O(n\Delta_{n}L_{n}^{5}),$
$\mathrm{tr}\bigl((M_{\sigma\alpha,n}^{\mathrm{o}})^{2}\bigr)$ $=O(n\Delta_{n}),$ $\mathrm{tr}\bigl((M_{H\alpha,n}^{\mathrm{o}})^{2}\bigr)$ $=O(n\Delta_{n}L_{n}^{2}),$ $\mathrm{tr}\bigl((M_{\alpha\alpha,n}^{\mathrm{o}})^{2}\bigr)$ $=O(n\Delta_{n}).$
Hence the pure $(\sigma,H)$ contributions are treated exactly as in mfBm, whereas the additional mixed terms satisfy
$(\delta_{\sigma,n}^{\mathrm{o}}\delta_{\alpha,n}^{\mathrm{o}})\left\lVert M_{\sigma\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-1}L_{n}^{-1/2}\bigr),$
$(\delta_{H,n}^{\mathrm{o}}\delta_{\alpha,n}^{\mathrm{o}})\left\lVert M_{H\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-1}L_{n}^{-1/2}\bigr),$
$(\delta_{\alpha,n}^{\mathrm{o}})^{2}\left\lVert M_{\alpha\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-1}\bigr),$
with the same orders for the corresponding squared Frobenius contributions. Consequently
$\left\lVert\mathcal{R}_{n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}=O\bigl((n\Delta_{n})^{-1}\bigr),\qquad\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)=O\bigl((n\Delta_{n})^{-1}\bigr)\to 0.$
Since equation 30 gives $\left\lVert B_{n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}\to 0$, the mfOU claim follows as well. ∎
5.3 Final proof of the LAN theorems
Proof of theorem 9.
Fix $h\in\mathbb{R}^{2}$ and write
$\Sigma_{n}:=\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}}),\qquad\Sigma_{n,h}:=\Sigma_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m},n}(h)),\qquad Z_{n}:=\Sigma_{n}^{-1/2}X_{n}^{\mathrm{m}}\sim\mathcal{N}(0,I_{n}).$
Define
$\mathcal{S}_{n}(h):=\Sigma_{n}^{-1/2}(\Sigma_{n,h}-\Sigma_{n})\Sigma_{n}^{-1/2}.$
The standard Gaussian likelihood identity gives
$\ell_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m},n}(h))-\ell_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})=-\frac{1}{2}\log\det\bigl(I_{n}+\mathcal{S}_{n}(h)\bigr)-\frac{1}{2}Z_{n}^{\top}\Bigl((I_{n}+\mathcal{S}_{n}(h))^{-1}-I_{n}\Bigr)Z_{n}.$ (32)
By proposition 15 and corollary 22, the sandwiched local covariance admits a second-order approximation with an additional cubic remainder that is already negligible, and in particular $\left\lVert\mathcal{S}_{n}(h)\right\rVert_{\mathrm{op}}\to 0$. Hence for $n$ large enough we may apply lemma 14 to obtain
$\ell_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m},n}(h))-\ell_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})$ $=\frac{1}{2}Q_{n}\bigl(\mathcal{S}_{n}(h)\bigr)-\frac{1}{4}\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)$ (33)
$\quad-\frac{1}{2}\Bigl(Z_{n}^{\top}\mathcal{S}_{n}(h)^{2}Z_{n}-\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)\Bigr)-\frac{1}{2}R_{\log}\bigl(\mathcal{S}_{n}(h)\bigr)$
$\quad-\frac{1}{2}Z_{n}^{\top}R_{\mathrm{inv}}\bigl(\mathcal{S}_{n}(h)\bigr)Z_{n}.$ (34)
Now write $\mathcal{S}_{n}(h)=B_{n}^{\mathrm{m}}(h)+\mathcal{R}_{n}^{\mathrm{m}}(h)$. Since
$\frac{1}{2}Q_{n}\bigl(B_{n}^{\mathrm{m}}(h)\bigr)=h^{\top}\Xi_{n}^{\mathrm{m}}$
and
$\mathrm{Var}\!\left(\frac{1}{2}Q_{n}\bigl(\mathcal{R}_{n}^{\mathrm{m}}(h)\bigr)\right)=\frac{1}{2}\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)\to 0,$
we have
$\frac{1}{2}Q_{n}\bigl(\mathcal{S}_{n}(h)\bigr)=h^{\top}\Xi_{n}^{\mathrm{m}}+o_{P_{\vartheta_{\mathrm{m}}}}(1).$ (35)
For the quadratic term,
$\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)=\mathrm{tr}\bigl(B_{n}^{\mathrm{m}}(h)^{2}\bigr)+2\mathrm{tr}\bigl(B_{n}^{\mathrm{m}}(h)\mathcal{R}_{n}^{\mathrm{m}}(h)\bigr)+\mathrm{tr}\bigl((\mathcal{R}_{n}^{\mathrm{m}}(h))^{2}\bigr).$
Because $\mathrm{tr}(B_{n}^{\mathrm{m}}(h)^{2})=O(1)$ by proposition 11, the Cauchy–Schwarz inequality and the local linearization result yield
$\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)=\mathrm{tr}\bigl(B_{n}^{\mathrm{m}}(h)^{2}\bigr)+o(1).$
Applying proposition 11 once more gives
$\frac{1}{4}\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)=\frac{1}{4}\mathrm{tr}\bigl(B_{n}^{\mathrm{m}}(h)^{2}\bigr)+o(1)=\frac{1}{2}h^{\top}\Gamma_{\mathrm{m}}^{\mathrm{crit}}h+o(1).$ (36)
Moreover,
$\mathrm{Var}\Bigl(Z_{n}^{\top}\mathcal{S}_{n}(h)^{2}Z_{n}-\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)\Bigr)=2\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{4}\bigr)\leq 2\left\lVert\mathcal{S}_{n}(h)\right\rVert_{\mathrm{op}}^{2}\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)\to 0,$
so
$Z_{n}^{\top}\mathcal{S}_{n}(h)^{2}Z_{n}-\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)=o_{P}(1).$ (37)
Next, lemma 14 and the previous bounds imply
$R_{\log}\bigl(\mathcal{S}_{n}(h)\bigr)=O\!\left(\left\lVert\mathcal{S}_{n}(h)\right\rVert_{\mathrm{op}}\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)\right)=o(1).$ (38)
Also,
$\mathrm{tr}\Bigl(R_{\mathrm{inv}}\bigl(\mathcal{S}_{n}(h)\bigr)^{2}\Bigr)\leq C\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{6}\bigr)\leq C\left\lVert\mathcal{S}_{n}(h)\right\rVert_{\mathrm{op}}^{4}\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)\to 0,$
while
$\left|\mathrm{tr}\Bigl(R_{\mathrm{inv}}\bigl(\mathcal{S}_{n}(h)\bigr)\Bigr)\right|\leq C\left\lVert\mathcal{S}_{n}(h)\right\rVert_{\mathrm{op}}\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)=o(1).$
Therefore centered Gaussian quadratic-form bounds give
$Z_{n}^{\top}R_{\mathrm{inv}}\bigl(\mathcal{S}_{n}(h)\bigr)Z_{n}=o_{P}(1).$ (39)
Combining (33)–(39) yields
$\ell_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m},n}(h))-\ell_{n}^{\mathrm{m}}(\vartheta_{\mathrm{m}})=h^{\top}\Xi_{n}^{\mathrm{m}}-\frac{1}{2}h^{\top}\Gamma_{\mathrm{m}}^{\mathrm{crit}}h+o_{P_{\vartheta_{\mathrm{m}}}}(1).$
Finally, theorem 6 gives $\Xi_{n}^{\mathrm{m}}\xrightarrow[]{d}\mathcal{N}(0,\Gamma_{\mathrm{m}}^{\mathrm{crit}})$, so the mfBm experiment is LAN at $\vartheta_{\mathrm{m}}$. ∎
Proof of theorem 10.
Fix $h\in\mathbb{R}^{3}$ and write
$\Sigma_{n}:=\Sigma_{n}^{\mathrm{o}}(\vartheta_{\mathrm{o}}),\qquad\Sigma_{n,h}:=\Sigma_{n}^{\mathrm{o}}(\vartheta_{\mathrm{o},n}(h)),\qquad Z_{n}:=\Sigma_{n}^{-1/2}X_{n}^{\mathrm{o}}\sim\mathcal{N}(0,I_{n}).$
Define $\mathcal{S}_{n}(h)=\Sigma_{n}^{-1/2}(\Sigma_{n,h}-\Sigma_{n})\Sigma_{n}^{-1/2}$. As in (32),
$\ell_{n}^{\mathrm{o}}(\vartheta_{\mathrm{o},n}(h))-\ell_{n}^{\mathrm{o}}(\vartheta_{\mathrm{o}})=-\frac{1}{2}\log\det\bigl(I_{n}+\mathcal{S}_{n}(h)\bigr)-\frac{1}{2}Z_{n}^{\top}\Bigl((I_{n}+\mathcal{S}_{n}(h))^{-1}-I_{n}\Bigr)Z_{n}.$
By proposition 15, corollary 22 and lemma 14,
$\ell_{n}^{\mathrm{o}}(\vartheta_{\mathrm{o},n}(h))-\ell_{n}^{\mathrm{o}}(\vartheta_{\mathrm{o}})$ $=\frac{1}{2}Q_{n}\bigl(\mathcal{S}_{n}(h)\bigr)-\frac{1}{4}\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)$
$\quad-\frac{1}{2}\Bigl(Z_{n}^{\top}\mathcal{S}_{n}(h)^{2}Z_{n}-\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)\Bigr)$
$\quad-\frac{1}{2}R_{\log}\bigl(\mathcal{S}_{n}(h)\bigr)-\frac{1}{2}Z_{n}^{\top}R_{\mathrm{inv}}\bigl(\mathcal{S}_{n}(h)\bigr)Z_{n}.$ (40)
Write $\mathcal{S}_{n}(h)=B_{n}^{\mathrm{o}}(h)+\mathcal{R}_{n}^{\mathrm{o}}(h)$. Since
$\frac{1}{2}Q_{n}\bigl(B_{n}^{\mathrm{o}}(h)\bigr)=h^{\top}\Xi_{n}^{\mathrm{o}},\qquad\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)\to 0,$
we obtain
$\frac{1}{2}Q_{n}\bigl(\mathcal{S}_{n}(h)\bigr)=h^{\top}\Xi_{n}^{\mathrm{o}}+o_{P_{\vartheta_{\mathrm{o}}}}(1).$ (41)
Furthermore,
$\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)=\mathrm{tr}\bigl(B_{n}^{\mathrm{o}}(h)^{2}\bigr)+o(1).$
Now proposition 12 gives
$\frac{1}{4}\mathrm{tr}\bigl(B_{n}^{\mathrm{o}}(h)^{2}\bigr)=\frac{1}{2}h^{\top}\Gamma_{\mathrm{o}}^{\mathrm{crit}}h+o(1),$
because the mixed traces with $A_{n}^{\mathrm{o}}$ are negligible under the critical normalization exactly as in the proof of theorem 7. Hence
$\frac{1}{4}\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)=\frac{1}{2}h^{\top}\Gamma_{\mathrm{o}}^{\mathrm{crit}}h+o(1).$ (42)
The higher-order terms are controlled exactly as in the mfBm case since $\left\lVert\mathcal{S}_{n}(h)\right\rVert_{\mathrm{op}}\to 0$ and $\mathrm{tr}(\mathcal{S}_{n}(h)^{2})=O(1)$:
$Z_{n}^{\top}\mathcal{S}_{n}(h)^{2}Z_{n}-\mathrm{tr}\bigl(\mathcal{S}_{n}(h)^{2}\bigr)=o_{P}(1),\qquad R_{\log}\bigl(\mathcal{S}_{n}(h)\bigr)=o(1),\qquad Z_{n}^{\top}R_{\mathrm{inv}}\bigl(\mathcal{S}_{n}(h)\bigr)Z_{n}=o_{P}(1).$
Combining these estimates with (5.3)–(42) yields
$\ell_{n}^{\mathrm{o}}(\vartheta_{\mathrm{o},n}(h))-\ell_{n}^{\mathrm{o}}(\vartheta_{\mathrm{o}})=h^{\top}\Xi_{n}^{\mathrm{o}}-\frac{1}{2}h^{\top}\Gamma_{\mathrm{o}}^{\mathrm{crit}}h+o_{P_{\vartheta_{\mathrm{o}}}}(1).$
Finally, theorem 7 implies $\Xi_{n}^{\mathrm{o}}\xrightarrow[]{d}\mathcal{N}(0,\Gamma_{\mathrm{o}}^{\mathrm{crit}})$, so the mfOU experiment is LAN at the critical point. ∎
Remark 16.
The limiting information matrices $\Gamma_{\mathrm{m}}^{\mathrm{crit}}$ and $\Gamma_{\mathrm{o}}^{\mathrm{crit}}$ are deterministic and non-singular. Hence the critical experiments are genuinely LAN, not merely LAMN. In particular, the critical point is resolved by a single triangular local reparametrization.
6 Score tests at the critical boundary
We now derive score-type tests from the critical LAN expansions established in theorems 9 and 10. Throughout, let
$z_{p}:=\Phi^{-1}(p)$
denote the $p$-quantile of the standard normal distribution.
Since the LAN expansions are centered at the boundary point $H=3/4$, the relevant tests are naturally calibrated under the critical constraint $H=3/4$. In the presence of nuisance parameters $\sigma$ and $\alpha$, the test for the $H$-direction is based on the efficient score, obtained by projecting the central sequence onto the orthocomplement of the nuisance tangent space.
(mfBm).
Let
$\Xi_{n}^{\mathrm{m}}=\begin{pmatrix}\Xi_{\sigma,n}^{\mathrm{m}}\\[3.00003pt] \Xi_{H,n}^{\mathrm{m}}\end{pmatrix}$
be the normalized score vector in theorem 9, and write
$\Gamma_{\mathrm{m}}^{\mathrm{crit}}=\begin{pmatrix}\Gamma_{\sigma\sigma}^{\mathrm{m}}&\Gamma_{\sigma H}^{\mathrm{m}}\\ \Gamma_{H\sigma}^{\mathrm{m}}&\Gamma_{HH}^{\mathrm{m}}\end{pmatrix}$.
Treating $\sigma$ as nuisance, the efficient score for $H$ is
$\Delta_{H,n}^{\mathrm{eff,m}}:=\Xi_{H,n}^{\mathrm{m}}-\Gamma_{H\sigma}^{\mathrm{m}}\bigl(\Gamma_{\sigma\sigma}^{\mathrm{m}}\bigr)^{-1}\Xi_{\sigma,n}^{\mathrm{m}}=\Xi_{H,n}^{\mathrm{m}}-\frac{\sigma}{2}\Xi_{\sigma,n}^{\mathrm{m}},$
and the corresponding efficient information is
$I_{H}^{\mathrm{eff,m}}:=\Gamma_{HH}^{\mathrm{m}}-\Gamma_{H\sigma}^{\mathrm{m}}\bigl(\Gamma_{\sigma\sigma}^{\mathrm{m}}\bigr)^{-1}\Gamma_{\sigma H}^{\mathrm{m}}=\frac{3\sigma^{4}}{64}.$
Accordingly, we consider the standardized statistic
$T_{n}^{\mathrm{m}}:=\frac{\Delta_{H,n}^{\mathrm{eff,m}}}{\sqrt{I_{H}^{\mathrm{eff,m}}}}=\frac{8}{\sqrt{3}\,\sigma^{2}}\left(\Xi_{H,n}^{\mathrm{m}}-\frac{\sigma}{2}\Xi_{\sigma,n}^{\mathrm{m}}\right).$
(mfOU).
Let
$\Xi_{n}^{\mathrm{o}}=\begin{pmatrix}\Xi_{\sigma,n}^{\mathrm{o}}\\[3.00003pt] \Xi_{H,n}^{\mathrm{o}}\\[3.00003pt] \Xi_{\alpha,n}^{\mathrm{o}}\end{pmatrix}$
be the normalized score vector in theorem 10. Write
$\eta:=(\sigma,\alpha),\qquad\Xi_{\eta,n}^{\mathrm{o}}:=\begin{pmatrix}\Xi_{\sigma,n}^{\mathrm{o}}\\[3.00003pt] \Xi_{\alpha,n}^{\mathrm{o}}\end{pmatrix}$,
and partition the critical information matrix as
$\Gamma_{\mathrm{o}}^{\mathrm{crit}}=\begin{pmatrix}\Gamma_{\eta\eta}^{\mathrm{o}}&\Gamma_{\eta H}^{\mathrm{o}}\\ \Gamma_{H\eta}^{\mathrm{o}}&\Gamma_{HH}^{\mathrm{o}}\end{pmatrix}$.
The efficient score for $H$ is then
$\Delta_{H,n}^{\mathrm{eff,o}}:=\Xi_{H,n}^{\mathrm{o}}-\Gamma_{H\eta}^{\mathrm{o}}\bigl(\Gamma_{\eta\eta}^{\mathrm{o}}\bigr)^{-1}\Xi_{\eta,n}^{\mathrm{o}}.$
Since (11) yields $\Gamma_{H\alpha}^{\mathrm{o}}=0$, this reduces to
$\Delta_{H,n}^{\mathrm{eff,o}}=\Xi_{H,n}^{\mathrm{o}}-\frac{\sigma}{2}\Xi_{\sigma,n}^{\mathrm{o}}.$
Likewise,
$I_{H}^{\mathrm{eff,o}}:=\Gamma_{HH}^{\mathrm{o}}-\Gamma_{H\eta}^{\mathrm{o}}\bigl(\Gamma_{\eta\eta}^{\mathrm{o}}\bigr)^{-1}\Gamma_{\eta H}^{\mathrm{o}}=\frac{3\sigma^{4}}{64},$
and we define
$T_{n}^{\mathrm{o}}:=\frac{\Delta_{H,n}^{\mathrm{eff,o}}}{\sqrt{I_{H}^{\mathrm{eff,o}}}}=\frac{8}{\sqrt{3}\,\sigma^{2}}\left(\Xi_{H,n}^{\mathrm{o}}-\frac{\sigma}{2}\Xi_{\sigma,n}^{\mathrm{o}}\right).$
Corollary 17.
Assume the critical LAN expansions in theorems 9 and 10.
  1. (i)
    Under the critical null hypothesis $H=3/4$,
    $T_{n}^{\mathrm{m}}\xrightarrow[]{d}\mathcal{N}(0,1),\qquad T_{n}^{\mathrm{o}}\xrightarrow[]{d}\mathcal{N}(0,1).$
  2. (ii)
    Under local alternatives with positive $H$-component, namely $h=(\cdot,h_{H})$ in the mfBm case and $h=(\cdot,h_{H},\cdot)$ in the mfOU case with $h_{H}>0$,
    $T_{n}^{\mathrm{m}}\xrightarrow[]{d}\mathcal{N}\!\bigl(h_{H}\sqrt{I_{H}^{\mathrm{eff,m}}},1\bigr),\qquad T_{n}^{\mathrm{o}}\xrightarrow[]{d}\mathcal{N}\!\bigl(h_{H}\sqrt{I_{H}^{\mathrm{eff,o}}},1\bigr).$
  3. (iii)
    Therefore, an asymptotic level-$\alpha$ boundary calibrated one-sided test of
    $H_{0}:\ H\leq\frac{3}{4}\qquad\text{vs}\qquad H_{1}:\ H>\frac{3}{4}$
    rejects for large values of the statistic:
    $T_{n}^{\mathrm{m}}>z_{1-\alpha}\quad\text{or}\quad T_{n}^{\mathrm{o}}>z_{1-\alpha},$
    according to the model under consideration.
Proof.
Under the critical null hypothesis $H=3/4$, the LAN central sequences in Theorems 9 and 10 satisfy
$\Xi_{n}^{\mathrm{m}}\xrightarrow[]{d}\mathcal{N}(0,\Gamma_{\mathrm{m}}^{\mathrm{crit}}),\qquad\Xi_{n}^{\mathrm{o}}\xrightarrow[]{d}\mathcal{N}(0,\Gamma_{\mathrm{o}}^{\mathrm{crit}}).$
We first treat the mfBm model. The nuisance parameter is $\sigma$. Hence the efficient $H$-score is the residual after projecting the $H$-component onto the nuisance direction
$\Delta_{H,n}^{\mathrm{eff,m}}:=\Xi^{\mathrm{m}}_{H,n}-\Gamma^{\mathrm{m}}_{H\sigma}\bigl(\Gamma^{\mathrm{m}}_{\sigma\sigma}\bigr)^{-1}\Xi^{\mathrm{m}}_{\sigma,n}.$
By the explicit entries of $\Gamma_{\mathrm{m}}^{\mathrm{crit}}$,
$\Gamma^{\mathrm{m}}_{H\sigma}\bigl(\Gamma^{\mathrm{m}}_{\sigma\sigma}\bigr)^{-1}=\frac{\sigma}{2},$
so
$\Delta_{H,n}^{\mathrm{eff,m}}=\Xi^{\mathrm{m}}_{H,n}-\frac{\sigma}{2}\Xi^{\mathrm{m}}_{\sigma,n}.$
Its asymptotic variance is the Schur complement
$I_{H}^{\mathrm{eff,m}}=\Gamma^{\mathrm{m}}_{HH}-\Gamma^{\mathrm{m}}_{H\sigma}\bigl(\Gamma^{\mathrm{m}}_{\sigma\sigma}\bigr)^{-1}\Gamma^{\mathrm{m}}_{\sigma H}.$
Therefore,
$T_{n}^{\mathrm{m}}=\frac{\Delta_{H,n}^{\mathrm{eff,m}}}{\sqrt{I_{H}^{\mathrm{eff,m}}}}\xrightarrow[]{d}\mathcal{N}(0,1).$
The mfOU case is analogous. Write
$\Xi_{n}^{\mathrm{o}}=\begin{pmatrix}\Xi^{\mathrm{o}}_{\eta,n}\\ \Xi^{\mathrm{o}}_{H,n}\end{pmatrix},\qquad\eta:=(\sigma,\alpha),\qquad\Gamma_{\mathrm{o}}^{\mathrm{crit}}=\begin{pmatrix}\Gamma^{\mathrm{o}}_{\eta\eta}&\Gamma^{\mathrm{o}}_{\eta H}\\ \Gamma^{\mathrm{o}}_{H\eta}&\Gamma^{\mathrm{o}}_{HH}\end{pmatrix}$.
Then
$\Delta_{H,n}^{\mathrm{eff,o}}:=\Xi^{\mathrm{o}}_{H,n}-\Gamma^{\mathrm{o}}_{H\eta}\bigl(\Gamma^{\mathrm{o}}_{\eta\eta}\bigr)^{-1}\Xi^{\mathrm{o}}_{\eta,n}.$
Since $\Gamma^{\mathrm{o}}_{H\alpha}=0$, this reduces to
$\Delta_{H,n}^{\mathrm{eff,o}}=\Xi^{\mathrm{o}}_{H,n}-\frac{\sigma}{2}\Xi^{\mathrm{o}}_{\sigma,n},$
and its variance is
$I_{H}^{\mathrm{eff,o}}=\Gamma^{\mathrm{o}}_{HH}-\Gamma^{\mathrm{o}}_{H\eta}\bigl(\Gamma^{\mathrm{o}}_{\eta\eta}\bigr)^{-1}\Gamma^{\mathrm{o}}_{\eta H}.$
Hence
$T_{n}^{\mathrm{o}}=\frac{\Delta_{H,n}^{\mathrm{eff,o}}}{\sqrt{I_{H}^{\mathrm{eff,o}}}}\xrightarrow[]{d}\mathcal{N}(0,1).$
This proves part (i).
For part (ii), under the local alternatives in Theorems 9 and 10, the LAN expansion holds with central sequence $\Xi_{n}^{\mathrm{m}}$ or $\Xi_{n}^{\mathrm{o}}$. By Le Cam’s third lemma ([5]), under a local alternative with $H$-component $h_{H}$, the efficient score acquires the mean shift
$h_{H}I_{H}^{\mathrm{eff,m}}\qquad\text{or}\qquad h_{H}I_{H}^{\mathrm{eff,o}},$
while its variance remains $I_{H}^{\mathrm{eff,m}}$ or $I_{H}^{\mathrm{eff,o}}$, respectively. After standardization, this gives
$T_{n}^{\mathrm{m}}\xrightarrow[]{d}\mathcal{N}\!\bigl(h_{H}\sqrt{I_{H}^{\mathrm{eff,m}}},\,1\bigr),\qquad T_{n}^{\mathrm{o}}\xrightarrow[]{d}\mathcal{N}\!\bigl(h_{H}\sqrt{I_{H}^{\mathrm{eff,o}}},\,1\bigr).$
Finally, part (iii) follows from part (i). Since the mean shift in part (ii) is positive when $h_{H}>0$, this is the natural right-tail rule for testing $H_{0}:H\leq 3/4$ against $H_{1}:H>3/4$. ∎
Corollary 18.
Under the assumptions of corollary 17, an asymptotic level-$\alpha$ two-sided test of
$H_{0}:\ H=\frac{3}{4}\qquad\text{vs}\qquad H_{1}:\ H\neq\frac{3}{4}$
rejects whenever
$|T_{n}^{\mathrm{m}}|>z_{1-\alpha/2}\quad\text{or}\quad|T_{n}^{\mathrm{o}}|>z_{1-\alpha/2},$
according to the model under consideration.
Practical implementation.
The preceding corollaries are stated for the oracle boundary statistics $T_{n}^{\bullet}$, that is, the statistics obtained when $\sigma$, $\alpha$ are treated as known. For numerical simulatuon, For numerical simulation, the true nuisance parameters $\sigma$ and $\alpha$ are replaced by restricted estimators computed under the constraint $H=3/4$.
  1. 1.
    Fix $H=3/4$ and estimate the nuisance parameters under this restriction. For mfBm this yields a restricted estimator $\widehat{\sigma}_{0}$, and for mfOU one analogously obtains $(\widehat{\sigma}_{0},\widehat{\alpha}_{0})$.
  2. 2.
    Evaluate the score components at the restricted estimates and at $H=3/4$.
  3. 3.
    Form the normalized score vector and project it onto the efficient $H$-direction by the formulas above.
  4. 4.
    Standardize by the corresponding plug-in efficient information and compare the resulting statistic with the Gaussian critical value.
This plug-in construction is the numerical counterpart of the boundary-calibrated score test used in the simulation study. Its full asymptotic justification under a fully composite null hypothesis for all parameters is left for future work.
Remark 19.
For implementation, the oracle statistics above may be replaced by plug-in versions obtained after estimating the nuisance parameters under the restricted constraint $H=3/4$.
Remark 20.
At the boundary point H=3/4, the opposite one-sided direction is handled by the same statistic with a left-tail rejection region.
$T^{\bullet}_{n}<z_{\alpha},\qquad\bullet\in\{\mathrm{m},\mathrm{o}\}.$
7 Simulation study
This section presents a first finite sample study of the plug-in one-sided critical score test. We examine how the statistic behaves when the nuisance parameter is estimated under the restriction $H=3/4$, and whether it can distinguish the critical case $H=3/4$ from a supercritical alternative $H>3/4$. The aim is numerical illustration rather than a formal asymptotic justification under the composite null hypothesis.
We only simulate mfBm here. For the mfOU model, the critical score structure in the $(\sigma,H)$-block is the same, while the $\alpha$-direction is asymptotically orthogonal at the leading order (see (11)). Thus the main problem already appears in the mfBm model and how the restricted nuisance fit enters the feasible $H$-score test. We therefore report only mfBm simulations here and leave the mfOU case for future work.
In every replication we impose the null hypothesis restriction $H=3/4$ and compute the restricted Gaussian MLE
$\widehat{\sigma}_{0}:=\arg\max_{\sigma>0}\ell_{n}^{\mathrm{m}}(\sigma,3/4).$
We then evaluate the mfBm score components at $(\widehat{\sigma}_{0},3/4)$ and form the plug-in efficient statistic
$\widehat{T}_{n}^{\mathrm{m}}=\frac{8}{\sqrt{3}\,\widehat{\sigma}_{0}^{2}}\left(\widehat{\Xi}_{H,n}^{\mathrm{m}}-\frac{\widehat{\sigma}_{0}}{2}\widehat{\Xi}_{\sigma,n}^{\mathrm{m}}\right),$
where
$\widehat{\Xi}_{\sigma,n}^{\mathrm{m}}=\frac{S_{\sigma,n}^{\mathrm{m}}(\widehat{\sigma}_{0},3/4)}{\sqrt{n\Delta_{n}L_{n}}},\qquad\widehat{\Xi}_{H,n}^{\mathrm{m}}=\frac{R_{H,n}^{\mathrm{m}}(\widehat{\sigma}_{0},3/4)}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}.$
The one-sided test rejects for large positive values of $\widehat{T}_{n}^{\mathrm{m}}$.
We present two numerical experiments.
Experiment I: Fixed design finite sample performance.
This experiment keeps the sampling design fixed and asks whether the statistic defined above already moves in the correct direction when the true Hurst index crosses from the critical point into the supercritical region. We take
$n=160,\qquad\Delta_{n}=10^{-2},\qquad\sigma=3,$
and use $200$ Monte Carlo replications in each regime. We compare
$H=3/4\qquad\text{and}\qquad H=0.90.$
The first is the critical null hypothesis, while the second is a clearly supercritical alternative. For each replication we compute $\widehat{\sigma}_{0}$, then $\widehat{T}_{n}^{\mathrm{m}}$, and finally the empirical rejection frequencies at the one-sided $10\%$ and $5\%$ Gaussian critical values.
Experiment II: Critical null hypothesis behavior along an asymptotic sequence.
The second experiment keeps the true model at the critical point
$H=3/4,\qquad\sigma=3,$
and examines how the plug-in fit and the resulting statistic behave along a simple asymptotic sequence. To approximate the LAN setting with $\Delta_{n}\downarrow 0$ and $n\Delta_{n}\to\infty$, we take
$\Delta_{n}=n^{-1/2},\qquad n\in\{64,128,256\},$
again with $200$ Monte Carlo replications for each $n$. The goal is descriptive: we track the numerical stability of $\widehat{\sigma}_{0}$ and compare the empirical behavior of $\widehat{T}_{n}^{\mathrm{m}}$ with its $\mathcal{N}(0,1)$ benchmark as the sample size increases.
Table 1: Experiment I: feasible one-sided mfBm test at a fixed design
[…]
$0.75$ 0.003-0.003 $0.532$ $0.025$ $0.005$ $3.034$ $0.377$
$0.90$ $1.461$ $1.726$ $0.445$ $0.380$ $1.519$ $0.456$
Table 2: Experiment II: critical-null hypothesis behavior of the plug-in mfBm statistic along an asymptotic sequence
[…]
$64$ $0.1250$ $0.006$ $0.751$ $0.045$ $2.948$ $0.358$
$128$ $0.0884$ $0.033$ $0.613$ $0.035$ $2.961$ $0.308$
$256$ $0.0625$ 0.047-0.047 $0.491$ $0.005$ $2.988$ $0.216$
The simulations show that the feasible mfBm statistic is numerically implementable once the restricted estimator $\widehat{\sigma}_{0}$ is obtained under $H=3/4$. Under the critical null hypothesis it is centered near zero but remains conservative in finite samples, suggesting slow convergence to the asymptotic normal limit. Under the supercritical alternative it shifts to the right and clearly improves rejection. Thus the current experiments support practical implementability and correct directional behavior, while a formal plug-in justification is left for future work.
Appendix A Critical low frequency reductions and explicit constants
This appendix records the one-dimensional reductions that lead to the critical constants in (8)–(9).For the mfBm model, we start from the exact spectral density given in (2). From it, we extract two key quantities $K_{\mathrm{m}}$ and $\beta_{\mathrm{m}}$ at $H=3/4$.
We then perform a low frequency rescaling by setting $\lambda=\Delta_{n}u$, which defines the critical window on which the spectral behavior simplifies. On this window, we derive the reduced symbols which are the functions of the scaled frequency $u$ that approximate the original spectral density and its derivatives.
Using these reduced symbols, we evaluate certain logarithmic integrals (see Appendix B). These integrals produce the three asymptotic orders $n\Delta_{n}L_{n}$, $n\Delta_{n}L_{n}^{2}$, and $n\Delta_{n}L_{n}^{3}$, which determine the scaling of the score covariances.
Finally, the last subsection of this appendix provides bounds for higher-order derivatives and for the remainder terms that appear in the local expansion. These bounds are essential for the LAN proof in this paper.
A.1 mfBm: critical window and reduced profiles
At $H=3/4$ the signal-to-noise ratio is
$\gamma_{n}:=\sigma^{2}\Delta_{n}^{2H-1}=\sigma^{2}\Delta_{n}^{1/2}.$
Using (5), for every fixed $A>0$ and uniformly on $1\leq|u|\leq A/\Delta_{n}$ we have
$\gamma_{n}f_{3/4}(\Delta_{n}u)=\sigma^{2}\Delta_{n}^{1/2}K_{\mathrm{m}}|\Delta_{n}u|^{-1/2}\bigl(1+r_{\mathrm{m}}(\Delta_{n}u)\bigr)=\eta_{\mathrm{m}}|u|^{-1/2}\bigl(1+o(1)\bigr),$ (43)
where $\eta_{\mathrm{m}}=\sigma^{2}K_{\mathrm{m}}$. Likewise,
$\gamma_{n}\,\partial_{H}f_{H}\big|_{H=3/4}(\Delta_{n}u)$ $=\eta_{\mathrm{m}}|u|^{-1/2}\Bigl(\beta_{\mathrm{m}}-2\log|\Delta_{n}u|+o(1)\Bigr)$
$=\eta_{\mathrm{m}}|u|^{-1/2}\Bigl(2L_{n}+\beta_{\mathrm{m}}-2\log|u|+o(1)\Bigr).$ (44)
Therefore the reduced covariance symbol
$a_{\mathrm{m},n}(u):=\frac{\gamma_{n}f_{3/4}(\Delta_{n}u)}{1+\gamma_{n}f_{3/4}(\Delta_{n}u)}$
converges locally uniformly on $\mathbb{R}\setminus\{0\}$ to
$w_{\mathrm{m}}(u)=\frac{\eta_{\mathrm{m}}|u|^{-1/2}}{1+\eta_{\mathrm{m}}|u|^{-1/2}}.$
Consequently the sandwiched score symbol for $C_{n}^{\mathrm{m}}$ is
$g_{\sigma}^{\mathrm{m}}(u)=\frac{2}{\sigma}w_{\mathrm{m}}(u),$
and, after removing the exact linear term $\sigma L_{n}\,S_{\sigma,n}^{\mathrm{m}}$, the reduced symbol associated with $D_{n}^{\mathrm{m}}$ is
$g_{H,n}^{\mathrm{m}}(u)=w_{\mathrm{m}}(u)\bigl(2L_{n}+\beta_{\mathrm{m}}-2\log|u|\bigr)+o\bigl(w_{\mathrm{m}}(u)\bigr),$
which is precisely the profile used in (20). In particular, for large $|u|$,
$w_{\mathrm{m}}(u)\sim\eta_{\mathrm{m}}|u|^{-1/2},\qquad g_{\sigma}^{\mathrm{m}}(u)\sim\frac{2\eta_{\mathrm{m}}}{\sigma}|u|^{-1/2},\qquad g_{H,n}^{\mathrm{m}}(u)\sim\eta_{\mathrm{m}}|u|^{-1/2}\bigl(2L_{n}+\beta_{\mathrm{m}}-2\log|u|\bigr).$ (45)
A.2 Explicit critical integrals for the mfBm block
Introduce the three basic logarithmic integrals
$J_{0}(L_{n})$ $:=2\int_{1}^{1/\Delta_{n}}u^{-1}\,\mathrm{d}u=2L_{n},$ (46)
$J_{1}(\beta,L_{n})$ $:=2\int_{1}^{1/\Delta_{n}}u^{-1}\bigl(2L_{n}+\beta-2\log u\bigr)\,\mathrm{d}u=2L_{n}^{2}+2\beta L_{n},$ (47)
$J_{2}(\beta,L_{n})$ $:=2\int_{1}^{1/\Delta_{n}}u^{-1}\bigl(2L_{n}+\beta-2\log u\bigr)^{2}\,\mathrm{d}u=\frac{8}{3}L_{n}^{3}+4\beta L_{n}^{2}+2\beta^{2}L_{n}.$ (48)
The bounded region $|u|\leq 1$ contributes only lower-order terms, so by (45),
$\frac{n\Delta_{n}}{2\pi}\int_{-1/\Delta_{n}}^{1/\Delta_{n}}\bigl(g_{\sigma}^{\mathrm{m}}(u)\bigr)^{2}\,\mathrm{d}u$ $\sim\frac{n\Delta_{n}}{2\pi}\frac{4\eta_{\mathrm{m}}^{2}}{\sigma^{2}}J_{0}(L_{n})=\frac{4\sigma^{2}K_{\mathrm{m}}^{2}}{\pi}\,n\Delta_{n}L_{n},$ (49)
$\frac{n\Delta_{n}}{2\pi}\int_{-1/\Delta_{n}}^{1/\Delta_{n}}g_{\sigma}^{\mathrm{m}}(u)g_{H,n}^{\mathrm{m}}(u)\,\mathrm{d}u$ $\sim\frac{n\Delta_{n}}{2\pi}\frac{2\eta_{\mathrm{m}}^{2}}{\sigma}J_{1}(\beta_{\mathrm{m}},L_{n})=\frac{2\sigma^{3}K_{\mathrm{m}}^{2}}{\pi}\,n\Delta_{n}L_{n}^{2},$ (50)
$\frac{n\Delta_{n}}{2\pi}\int_{-1/\Delta_{n}}^{1/\Delta_{n}}\bigl(g_{H,n}^{\mathrm{m}}(u)\bigr)^{2}\,\mathrm{d}u$ $\sim\frac{n\Delta_{n}}{2\pi}\eta_{\mathrm{m}}^{2}J_{2}(\beta_{\mathrm{m}},L_{n})=\frac{4\sigma^{4}K_{\mathrm{m}}^{2}}{3\pi}\,n\Delta_{n}L_{n}^{3}.$ (51)
These are exactly the constants in (8); after substituting $K_{\mathrm{m}}=3\sqrt{2\pi}/8$, the entries reduce to the explicit values $9\sigma^{2}/8$, $9\sigma^{3}/16$, and $3\sigma^{4}/8$. The point is that the leading term of $J_{1}$ is $2L_{n}^{2}$ and the leading term of $J_{2}$ is $(8/3)L_{n}^{3}$, so one obtains the critical covariance orders without any further projection.
A.3 mfOU: critical window and the drift profile
For the stationary mfOU model, write the sampled spectral density near the origin in the form
$f_{\Delta}^{\mathrm{o}}(\lambda;\sigma,3/4,\alpha)=K_{\mathrm{o}}|\lambda|^{-1/2}\bigl(1+r_{\mathrm{o},\Delta}(\lambda)\bigr),$
where $r_{\mathrm{o},\Delta}(\lambda)\to 0$ on low frequency windows. Then the same calculation as in (43)–(44) gives, uniformly on $1\leq|u|\leq A/\Delta_{n}$,
$\gamma_{n}f_{\Delta}^{\mathrm{o}}(\Delta_{n}u;\vartheta_{\mathrm{o}})$ $=\eta_{\mathrm{o}}|u|^{-1/2}\bigl(1+o(1)\bigr),$ (52)
$\gamma_{n}\,\partial_{H}f_{\Delta}^{\mathrm{o}}(\Delta_{n}u;\vartheta_{\mathrm{o}})$ $=\eta_{\mathrm{o}}|u|^{-1/2}\bigl(2L_{n}+\beta_{\mathrm{o}}-2\log|u|+o(1)\bigr),$ (53)
with $\eta_{\mathrm{o}}=\sigma^{2}K_{\mathrm{o}}$. Hence the $(\sigma,H)$-block has exactly the same critical reduced profiles as in the mfBm case, after replacing $(\eta_{\mathrm{m}},\beta_{\mathrm{m}})$ by $(\eta_{\mathrm{o}},\beta_{\mathrm{o}})$. In particular,
$\frac{n\Delta_{n}}{2\pi}\int_{-1/\Delta_{n}}^{1/\Delta_{n}}\bigl(g_{\sigma}^{\mathrm{o}}(u)\bigr)^{2}\,\mathrm{d}u$ $\sim\frac{4\sigma^{2}K_{\mathrm{o}}^{2}}{\pi}\,n\Delta_{n}L_{n},$
$\frac{n\Delta_{n}}{2\pi}\int_{-1/\Delta_{n}}^{1/\Delta_{n}}g_{\sigma}^{\mathrm{o}}(u)g_{H,n}^{\mathrm{o}}(u)\,\mathrm{d}u$ $\sim\frac{2\sigma^{3}K_{\mathrm{o}}^{2}}{\pi}\,n\Delta_{n}L_{n}^{2},$
$\frac{n\Delta_{n}}{2\pi}\int_{-1/\Delta_{n}}^{1/\Delta_{n}}\bigl(g_{H,n}^{\mathrm{o}}(u)\bigr)^{2}\,\mathrm{d}u$ $\sim\frac{4\sigma^{4}K_{\mathrm{o}}^{2}}{3\pi}\,n\Delta_{n}L_{n}^{3}.$ (54)
For the drift parameter, the reduced profile is
$g_{\alpha}^{\mathrm{o}}(u)=-\frac{2\alpha}{\alpha^{2}+u^{2}},$
so
$\int_{\mathbb{R}}\bigl(g_{\alpha}^{\mathrm{o}}(u)\bigr)^{2}\,\mathrm{d}u=4\alpha^{2}\int_{\mathbb{R}}\frac{\mathrm{d}u}{(\alpha^{2}+u^{2})^{2}}=\frac{2\pi}{\alpha}.$ (55)
Thus
$\frac{n\Delta_{n}}{2\pi}\int_{\mathbb{R}}\bigl(g_{\alpha}^{\mathrm{o}}(u)\bigr)^{2}\,\mathrm{d}u\sim\frac{1}{\alpha}\,n\Delta_{n},$
which yields the fully explicit value $\Xi_{\alpha\alpha}^{\mathrm{o}}=\alpha^{-1}$. Moreover, using the tails from (52)–(53),
$|g_{\sigma}^{\mathrm{o}}(u)g_{\alpha}^{\mathrm{o}}(u)|\lesssim\frac{|u|^{-1/2}}{1+u^{2}},\qquad|g_{H,n}^{\mathrm{o}}(u)g_{\alpha}^{\mathrm{o}}(u)|\lesssim\frac{|u|^{-1/2}(L_{n}+|\log|u||)}{1+u^{2}}.$
Both dominating functions are integrable on $\mathbb{R}$; the second one has integral $O(L_{n})$ because only the explicit prefactor $L_{n}$ grows with $n$. Therefore
$\frac{n\Delta_{n}}{2\pi}\int_{\mathbb{R}}|g_{\sigma}^{\mathrm{o}}(u)g_{\alpha}^{\mathrm{o}}(u)|\,\mathrm{d}u=O(n\Delta_{n}),\qquad\frac{n\Delta_{n}}{2\pi}\int_{\mathbb{R}}|g_{H,n}^{\mathrm{o}}(u)g_{\alpha}^{\mathrm{o}}(u)|\,\mathrm{d}u=O(n\Delta_{n}L_{n}),$
which is exactly the mixed-trace order stated in (31).
A.4 Higher-order critical profiles and LAN remainders
The LAN proof itself only uses the first-order local perturbation $B_{n}^{\bullet}(h)$ and the quadratic information term, but the critical window calculus in fact controls one order deeper. The point of this subsection is to record those higher-order bounds in a form that can be reused later.
Proposition 21.
For either model $\bullet\in\{\mathrm{m},\mathrm{o}\}$, let
$N_{\beta,n}^{\bullet}(\theta):=\Sigma_{n}^{\bullet}(\vartheta_{\bullet})^{-1/2}\partial_{\theta}^{\beta}\Sigma_{n}^{\bullet}(\theta)\Sigma_{n}^{\bullet}(\vartheta_{\bullet})^{-1/2},\qquad|\beta|=2,3,$
and let $m(\beta)$ denote the number of $H$-derivatives appearing in $\beta$. Then, uniformly on shrinking neighborhoods of the critical point,
$\left\lVert N_{\beta,n}^{\bullet}\right\rVert_{\mathrm{op}}=O\bigl(L_{n}^{m(\beta)}\bigr),\qquad\mathrm{tr}\Bigl(\bigl(N_{\beta,n}^{\bullet}\bigr)^{2}\Bigr)=O\bigl(n\Delta_{n}L_{n}^{2m(\beta)+1}\bigr),\qquad|\beta|=2,3.$ (56)
In particular, every additional $H$-derivative costs exactly one extra logarithmic factor at the operator level and two at the trace level.
Proof.
At the critical point, every differentiation with respect to $H$ inserts one extra logarithmic factor into the reduced low frequency symbol, whereas $\sigma$- and $\alpha$-derivatives only modify smooth prefactors. Concretely, on the critical window $\lambda=\Delta_{n}u$ one has for mfBm
$\gamma_{n}\,\partial_{H}^{k}f_{H}\big|_{H=3/4}(\Delta_{n}u)=\eta_{\mathrm{m}}|u|^{-1/2}\Bigl(P_{k,\mathrm{m}}(2L_{n}-2\log|u|)+o\bigl((1+L_{n}+|\log|u||)^{k}\bigr)\Bigr),\qquad k=1,2,3,$
where $P_{k,\mathrm{m}}$ is a polynomial of degree $k$ with leading term $x^{k}$. The same description holds for the sampled mfOU symbol, with polynomials $P_{k,\mathrm{o}}$ of the same degree and different lower-order coefficients. Hence every squared reduced symbol contributing to $N_{\beta,n}^{\bullet}$ is dominated by
$|u|^{-1}\bigl(1+L_{n}+|\log|u||\bigr)^{2m(\beta)}.$
The Toeplitz operator norm is governed by the $L^{\infty}$ size of the reduced symbol, while the Frobenius norm is governed by its $L^{2}$ mass. Therefore it is enough to record the elementary logarithmic integral
$\int_{1}^{1/\Delta_{n}}u^{-1}\bigl(1+L_{n}+|\log u|\bigr)^{2m}\,\mathrm{d}u=O\bigl(L_{n}^{2m+1}\bigr),\qquad m\in\{0,1,2,3\},$ (57)
which is immediate after the change of variables $v=\log u$. The bound (56) follows by the same Toeplitz trace reduction as in the proof of the first-order score asymptotics. ∎
Corollary 22.
Write the sandwiched local covariance expansion one order deeper as
$\mathcal{S}_{n}^{\bullet}(h)=B_{n}^{\bullet}(h)+\mathcal{R}_{2,n}^{\bullet}(h)+\mathcal{R}_{3,n}^{\bullet}(h),$
where $\mathcal{R}_{2,n}^{\bullet}(h)$ is quadratic and $\mathcal{R}_{3,n}^{\bullet}(h)$ is cubic in the local shift. Then, for every fixed $h$,
$\left\lVert\mathcal{R}_{2,n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{2,n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$ (58)
$\left\lVert\mathcal{R}_{2,n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-1}\bigr),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{2,n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)$ $=O\bigl((n\Delta_{n})^{-1}\bigr),$ (59)
$\left\lVert\mathcal{R}_{3,n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-3/2}L_{n}^{-3/2}\bigr),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{3,n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)$ $=O\bigl((n\Delta_{n})^{-2}L_{n}^{-2}\bigr),$ (60)
$\left\lVert\mathcal{R}_{3,n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-3/2}\bigr),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{3,n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)$ $=O\bigl((n\Delta_{n})^{-2}\bigr).$ (61)
Consequently the cubic local covariance contribution is already negligible before the matrix Taylor expansion for the log-determinant and the inverse is applied.
Proof.
The local shifts satisfy
$\delta_{\sigma,n}=O\bigl((n\Delta_{n}L_{n})^{-1/2}\bigr),\qquad\delta_{H,n}=O\bigl((n\Delta_{n})^{-1/2}L_{n}^{-3/2}\bigr),\qquad\delta_{\alpha,n}=O\bigl((n\Delta_{n})^{-1/2}\bigr).$
For mfBm, the quadratic remainder contains only the three terms
$(\delta_{\sigma,n})^{2}N_{(\sigma\sigma),n}^{\mathrm{m}},\qquad\delta_{\sigma,n}\delta_{H,n}N_{(\sigma H),n}^{\mathrm{m}},\qquad(\delta_{H,n})^{2}N_{(HH),n}^{\mathrm{m}},$
and proposition 21 shows that each of them is of operator order $(n\Delta_{n}L_{n})^{-1}$ and squared-trace order $(n\Delta_{n}L_{n})^{-1}$. This yields (58). For mfOU the additional $(\sigma,\alpha)$, $(H,\alpha)$ and $(\alpha,\alpha)$ terms are even smaller, except for $(\alpha,\alpha)$ whose size is exactly $(n\Delta_{n})^{-1}$; hence (59) follows.
For the cubic remainder, every monomial contains three local shifts and a third derivative. In mfBm the worst case is any monomial with total $m(\beta)=3$ after multiplying by $(\delta_{H,n})^{3}$, and all such contributions have the same size
$(n\Delta_{n})^{-3/2}L_{n}^{-9/2}\cdot L_{n}^{3}=(n\Delta_{n})^{-3/2}L_{n}^{-3/2}$
at the operator level. Squaring and multiplying by the trace order $n\Delta_{n}L_{n}^{2\cdot 3+1}$ gives (60). The mfOU bound is identical, except that the pure $\alpha$-terms dominate and remove the negative logarithmic power, which yields (61). Therefore all terms beyond the quadratic LAN term are absorbed into the same $o_{P}(1)$ remainder. ∎
Appendix B Elementary logarithmic integrals
For quick reference, the three elementary logarithmic identities used repeatedly throughout the critical analysis are
$\int_{1}^{1/\Delta_{n}}u^{-1}\,\mathrm{d}u$ $=L_{n},$
$\int_{1}^{1/\Delta_{n}}u^{-1}\log u\,\mathrm{d}u$ $=\frac{1}{2}L_{n}^{2},$
$\int_{1}^{1/\Delta_{n}}u^{-1}\log^{2}u\,\mathrm{d}u$ $=\frac{1}{3}L_{n}^{3}.$
They are the sole source of the critical orders $n\Delta_{n}L_{n}$, $n\Delta_{n}L_{n}^{2}$, and $n\Delta_{n}L_{n}^{3}$ appearing in the score covariance and in the LAN quadratic term.